### v1

# An order-theoretic characterization of JB-algebras

Abstract
We give an order-theoretic characterization of the JB-algebras among the complete order unit spaces in terms of the existence of an order-anti-automorphism of the interior of the cone that is homogeneous of degree $-1$. More geometrically, we characterize JB-algebras as those complete order unit spaces for which the interior of the cone is a symmetric Banach–Finsler manifold under Thompson’s metric. Furthermore, we show that two order unit spaces are isomorphic if there exists a gauge-reversing bijection between them, thus answering a question raised by Noll–Schäfer. These results have previously been established for finite-dimensional resp. reflexive order unit spaces by Walsh and Lemmens–R.–Wortel.
Keywords: Jordan algebras, JB-algebras, order unit spaces, order-anti-isomorphisms, gauge-reversing bijections, Thompson’s metric, symmetric Banach–Finsler manifolds
Subject Classification: 17C50, 58B20, 46B40, 17C36
1 Introduction
Jordan algebras form a fecund class of nonassociative real algebras. Originating in the work of P. Jordan, J. von Neumann, and E. Wigner [JvNW34] on the foundations of quantum mechanics in the 1930s, Jordan algebras have since been connected to a vast range of fields including optimization (cf.[AG03] [Fay97]), statistics (cf. [Mal94]), complex and differential geometry (cf. [Koe99], [Ber00]), Lie theory (cf. [FF77], [Jac71]), and the study of symmetric spaces (cf. [Loo69], [Kau02], [Upm85], [Sat80]). More specifically, we consider a unital Jordan algebra $A$ which is simultaneously a Banach space for a norm satisfying $\lVert xy\rVert\leq\lVert x\rVert\lVert y\lVert$ and $\lVert x\rVert^{2}\leq\lVert x^{2}+y^{2}\rVert$ for all $x,y\in A$; such pairs $(A,\lVert\cdot\rVert)$ are known as JB-algebras. The purpose of this paper is to establish a conjectured correspondence between JB-algebras and certain so-called order unit spaces.
An order unit space $(V,V_{+},v)$ is a partially ordered vector space $V$, with positive cone $V_{+}$, together with a choice of archimedean order unit $v\in V$, see Section 2.4. One defines intrinsically a norm $\lVert\cdot\rVert_{v}$ on $V$ deemed the order unit norm. The $\lVert\cdot\rVert_{v}$-interior of $V_{+}$ is the open cone
$V^{\circ}_{+}:=\{x\in V:x\geq\delta v\text{ for some }\delta>0\}.$
If $A$ is a JB-algebra with unit element $e$ and one sets $A_{+}:=\{x^{2}:x\in A\}$, then the triple $(A,A_{+},e)$ is an example of an order unit space. Conversely, the question how to characterize those order unit spaces that arise in this way from a JB-algebra has received sustained attention. We provide an answer based on the following observation. The $\lVert\cdot\rVert_{e}$-interior $A^{\circ}_{+}$ of $A_{+}$ consists precisely of the squares having a Jordan inverse. The Jordan inverse map $i\colon A^{\circ}_{+}\to A^{\circ}_{+}$, $i(x)=x^{-1},$ is a bijection having the properties of being an order-anti-isomorphism and homogeneous of degree $-1$, that is $x\leq y\iff i(x)\geq i(y)$ and $i(\lambda x)=\lambda^{-1}i(x)$ for all $x,y\in A^{\circ}_{+}$ and $\lambda>0$. This observation motivates the following general definition.
Let $(V,V_{+},v)$ be an order unit space. A bijection $\Phi:V^{\circ}_{+}\to V^{\circ}_{+}$ is called gauge-reversing if it is an order-anti-isomorphism and is homogeneous of degree $-1$. (For an alternate definition in terms of so-called gauge functions, see Section 2.6.) Our main result is the following order-theoretic characterization of JB-algebras among the order unit spaces, which resolves [LvG18, Conjecture 2].
Theorem 1.1.
Let $(V,V_{+},v)$ be a complete order unit space. Then there exists a structure of Jordan algebra on $V$ such that $(V,\lVert\cdot\rVert_{v})$ is a JB-algebra with unit element $v$ and cone of squares $V_{+}$ if and only if there exists a gauge-reversing bijection $V^{\circ}_{+}\to V^{\circ}_{+}$.
We will prove this theorem as Theorem 5.33.
Let us now compare Theorem 1.1 with other known characterizations of JB-algebras. This line of research has its origins in the Koecher-Vinberg theorem ([Koe57] and [Vin60]) asserting that a cone in a finite-dimensional real vector space corresponds to a JB-algebra (or what is equivalent, a Euclidean Jordan algebra [FK94, Ch. III, Sec. 1; Ch. VIII, Sec. 4], or a formally real Jordan algebra [HOS84, Cor. 3.1.7 and 3.3.8]) if and only if the cone is symmetric, i.e. homogeneous and self-dual for some inner product. This result has been extended to characterize unital JH-algebras [Chu17], which are precisely the reflexive JB-algebras (see [LRW25, Thm. 2.8]). A characterization of general, not necessarily reflexive, JB-algebras, however, cannot involve the existence of an inner product and has been obtained by Walsh.
Walsh has shown that an open cone in a finite-dimensional vector space is symmetric if and only if it admits a gauge-reversing self-bijection [Wal18]. Combined with the Koecher-Vinberg theorem, Walsh’s result proves Theorem 1.1 in case $V$ is finite-dimensional. This result has been extended to characterize spin factors [LRvI17] and JH-algebras [LRW25], relying on the aforementioned extension of the Koecher-Vinberg theorem to infinite dimension by Chu. However, in the general case a different method is required to construct the algebra multiplication.
This method is provided by the foundational work of Springer on $j$-structures [Spr73]. He defined a $j$-structure, roughly speaking, to be a birational involution on a finite dimensional vector space $V$ that is homogeneous of degree $-1$ and satisfies the weak Hua identity
$j(v+x)+j(v+j(x))=v=j(v).$
Springer showed that $j$-structures are equivalent to finite-dimensional Jordan algebras in characteristic $\neq 2$. McCrimmon adopted the definition of a $j$-structure to what he coined an $h$-structure, which is required to satisfy a stronger form of Hua’s identity [McC77, Sec. 3]. It is this notion of an $h$-structure which we will implicitly use, although we will not use this term.
JB-algebras have also been studied from the perspective of symmetric spaces. In [KU77] and [BKU78] it has been shown that an open cone $C$ in a Banach space has a JB-algebra structure (in an equivalent norm) if and only if the tube domain $T=V\oplus\mathrm{i}C\subset V^{\mathbb{C}}$ over $C$ admits a symmetry $s$, i.e. a holomorphic involution $s\colon T\to T$ with an isolated fixed point, equivalently, if and only $T$ is biholomorphic to a bounded symmetric domain.
One obtains a so-called Finsler metric $\nu\colon C\times C\to[0,\infty)$ on the open cone $C=V_{+}^{\circ}$ of an order unit space $V$ from a suitable assignment to each point in $C$ of a norm on $V$ which is equivalent to $\lVert\cdot\rVert_{v}$. An example of an $\operatorname{Aut}(C)$-invariant Finsler metric on $C$ is Thompson’s metric $d_{T}\colon C\times C\to[0,\infty)$ defined by
$d_{T}(x,y):=\log\inf\{\lambda\geq 1:\lambda^{-1}y\leq x\leq\lambda y\}.$
A $\nu$-symmetry of $C$ at a point $p\in C$ is an involutive $\nu$-isometry having $p$ as an isolated fixed point. We say that $(C,\nu)$ is a symmetric Banach–Finsler manifold if there exists a $\nu$-symmetry around each point of $C$. It has been shown in [LRW25, Thm. 3.5] that any $d_{T}$-symmetry is a gauge-reversing bijection.
It is known that the open cone of each JB-algebra is a symmetric Banach–Finsler manifold. In fact, Chu has proved in [Chu21] that $V$ corresponds to a JB-algebra if and only if there exist an $\operatorname{Aut}(C)$-invariant Finsler metric $\nu$ and analytic $\nu$-symmetries turning $(C,\nu)$ into a symmetric Banach–Finsler manifold. Without assuming any continuity or smoothness of the symmetries, we provide a new proof of Chu’s result in case $\nu=d_{T}$, thus answering in the affirmative [LvG18, Conjecture 1]. The following theorem follows from the preceding discussion and Theorem 1.1.
Theorem 1.2.
Let $V$ be a complete order unit space and let $C=V_{+}^{\circ}$ be its open cone. Then the following are equivalent:
  1. (1)
    there exists a $d_{T}$-symmetry $S_{p}:C\to C$ at some point $p\in C$;
  2. (2)
    we have that $(C,d_{T})$ is a symmetric Banach–Finsler manifold;
  3. (3)
    there exists a JB-algebra structure on $(V,\lVert\cdot\rVert_{v})$ with cone of squares $V_{+}$.
We will prove this theorem as Theorem 5.35.
An overview of the proof of Theorem 1.1 will now be given. Throughout this paper, we work in the following, more general, context. We assume $(V,V_{+},v)$ and $(W,W_{+},w)$ are two complete order unit spaces, with open cones $C=V_{+}^{\circ}$ and $D=W_{+}^{\circ}$, and that $\Phi\colon C\to D$ is a gauge-reversing bijection.
In Section 3 we define cones $C_{usc}$ and $C_{lsc}$ consisting of upper respectively lower semicontinuous, affine functions on the state space of $V$, which are obtained by adjoining to $C$ all limits of decreasing respectively increasing nets in $C$. The map $\Phi\colon C\to D$ is extended to a gauge-reversing bijection $\Phi_{usc}\colon C_{usc}\to D_{lsc}$. We establish that the cone $C_{usc}$ is strongly atomic, i.e. each element in $C_{usc}$ is the supremum of the extremal vectors that it dominates. In fact, we show that under $\Phi_{usc}$ the extremal vectors of $C_{usc}$ correspond to the pure states on $W$, whose ample existence is guaranteed by the Krein–Milman theorem.
In Section 4 several consequences of strong atomicity of $C_{usc}$ are deduced by scrutinizing the ordering. First, we show that $\Phi$ is Gateaux-differentiable in all positive directions in Theorem 4.7. Second, we establish a strong form of Hua’s identity in Theorem 4.10. Third, we reprove [NS77, Theorem B] asserting that a gauge-preserving bijection $\Psi\colon C\to D$ extends to a linear isomorphism from $V$ onto $W$ in the special case that $C_{usc}$ be strongly atomic in Theorem 4.9.
In Section 5 we prove that $\Phi$ is Fréchet-differentiable and that for each point $x\in C$ the negative of the derivative at $x$ defines an isomorphism $-D\Phi(x)\colon(V,V_{+},x)\to(W,W_{+},\Phi(x))$ of order unit spaces. We show that $S_{x}:=-[D\Phi(x)]^{-1}\circ\Phi$ is a $d_{T}$-symmetry of $C$ at $x$. Specifically, we let $j:=S_{v}$ be the prospective inverse map of the Jordan algebra structure on $V$ with identity element $v$. We define the quadratic representation of an element $x\in C$ by $P(x):=-[Dj(x)]^{-1}$ and show that $P$ extends to a quadratic polynomial map $P\colon V\to L(V)$. As a consequence we prove that every gauge-reversing bijection $\Phi\colon C\to D$ is bianalytic. The map $P$ is shown to satisfy an alternative axiomatization of JB-algebras in terms of the quadratic representation (see Definition 2.41), finishing the proof of Theorem 1.1.
As a byproduct of the proof, we answer in the affirmative a question raised by Noll and Schäfer [NS77] whether two possibly non-complete order unit spaces connected by a gauge-reversing bijection are necessarily isomorphic.
Theorem 1.3.
If $V$ and $W$ are order unit spaces such that there exists a gauge-reversing bijection $\Phi\colon C\to D$, then $V$ and $W$ are isomorphic.
Once more, in case $V$ and $W$ are finite-dimensional this was shown by Walsh [Wal18]. We prove Theorem 1.3 in Section 5.8 as Theorem 5.38, by showing that a gauge-reversing bijection $\Phi\colon C\to D$ extends to a gauge-reversing bijection $\hat{\Phi}\colon\hat{C}\to\hat{D}$, where $\hat{C}$ and $\hat{D}$ are the open cones of the norm completions of $V$ resp. $W$. We have also included a version of Theorem 1.1 for non-complete order unit spaces in that section.
The following section covers preliminary concepts and results concerning differentiable and analytic maps, order unit spaces, and JB-algebras.
2 Preliminaries
2.1 Differentiable maps
In this section, we recall the definition of Gateaux- and Fréchet-differentiable maps, and prove a chain rule for the Gateaux-derivative.
Definition 2.1.
Let $(X,\lVert\cdot\rVert_{X})$ and $(Y,\lVert\cdot\rVert_{Y})$ be Banach spaces. Let $F\subset X$ be a subset and $f\colon F\to Y$ a function.
  • $(1)$
    Given a point $x\in F$ and a vector $p\in X$, we say that $f$ is Gateaux-differentiable at $x$ in the direction $p$ if for sufficiently small $\mu>0$ we have $x+\mu p\in F$, and there is a vector $q\in Y$ such that
    $\lim_{\mu\downarrow 0}\mu^{-1}\lVert f(x+\mu p)-f(x)-\mu q\rVert_{Y}=0.$
    Plainly, the vector $q$ is unique if it exists, and called the Gateaux-derivative of $f$ at $x$ in the direction $p$. We emphasize that, due to the right limit in the definition, our notion of Gateaux-derivative is one-sided. Nonetheless, if $\lambda\geq 0$, then $\lambda q$ is the Gateaux-derivative of $f$ at $x$ in the direction $\lambda p$.
  • $(2)$
    Let $x$ belong to the interior of $F$. We say that $f$ is Fréchet-differentiable at $x$ if there exists a continuous linear map $A\colon X\to Y$ such that
    $\lim_{h\to 0}\lVert h\rVert_{X}^{-1}\lVert f(x+h)-f(x)-A(h)\rVert_{Y}=0.$
    Again, the linear map $A$ is unique if it exists. We call $A$ the Fréchet-derivative of $f$ at $x$ and denote it $Df(x)$. Note that if $f$ is Fréchet-differentiable at $x$, then $f$ is continuous at $x$.
We will use the following well-known chain rule for Fréchet- and Gateaux-derivatives.
Lemma 2.2.
Let $X,Y,Z$ be Banach spaces. Let $f\colon F\to G$ and $g\colon G\to Z$ be functions, defined on subsets $F\subset X$ and $G\subset Y$. Let $x\in F$ and set $y:=f(x)\in G$.
  • $(1)$
    If $f$ is Fréchet-differentiable at $x$, and $g$ is Fréchet-differentiable at $y$, then $g\circ f$ is Fréchet-differentiable at $x$ with Fréchet-derivative given by the chain rule
    $D(g\circ f)(x)=Dg(y)\circ Df(x).$
  • $(2)$
    Let $p\in X$, $q\in Y$ and $r\in Z$ be vectors such that $q$ is the Gateaux-derivative of $f$ at $x$ in the direction $p$, and $r$ is the Gateaux-derivative of $g$ at $y$ in the direction $q$. Assume additionally that $g$ is Lipschitz on a neighbourhood of $y$ in $G$. Then $r$ is the Gateaux derivative of $g\circ f$ at $x$ in the direction $p$.
Proof.
Part (1) is standard, so we only show part (2).
Since $g$ is Lipschitz on a neighborhood of $y$ in $G$, there exist $\delta>0$ and $L>0$ such that $\lVert g(y_{1})-g(y_{2})\rVert_{Z}\leq L\lVert y_{1}-y_{2}\rVert_{Y}$ whenever $y_{i}\in G$ with $\lVert y_{i}-y\rVert_{Y}<\delta$ for $i=1,2$. For $\mu>0$ sufficiently small we have $\lVert\mu q\rVert_{Y}<\delta/2$ and $\lVert f(x+\mu p)-f(x)-\mu q\lVert_{Y}<\delta/2$, hence $\lVert f(x+\mu p)-y\lVert<\delta$ and $\lVert f(x)+\mu q-y\rVert<\delta$. Since $g$ is $L$-Lipschitz on $\{y^{\prime}\in G:\lVert y^{\prime}-y\rVert_{Y}<\delta\}$, for sufficiently small $\mu>0$ we have
$\lVert g(f(x+\mu p))-g(f(x)+\mu q)\rVert_{Z}\leq L\lVert f(x+\mu p)-f(x)-\mu q\rVert_{Y}$
and hence by the triangle inequality
$\lVert(g\circ f)(x+\mu p)-(g\circ f)(x)-\mu r\rVert_{Z}$ $\leq L\lVert f(x+\mu p)-f(x)-\mu q\rVert_{Y}$
$\hskip 8.2511pt+\lVert g(f(x)+\mu q)-g(f(x))-\mu r\rVert_{Z}.$
As $q$ is the Gateaux-derivative of $f$ at $x$ in the direction $p$ and $r$ is the Gateaux-derivative of $g$ at $y$ in the direction $q$, it follows that $r$ is the Gateaux-derivative of $g\circ f$ at $x$ in the direction $p$. ∎
2.2 Analytic maps
Let $(X,\lVert\cdot\rVert_{X})$ and $(Y,\lVert\cdot\rVert_{Y})$ be Banach spaces. The norm of a multilinear map $\phi\colon X^{k}\to Y$ is given by
$\lVert\phi\rVert:=\sup\{\lVert\phi(x_{1},x_{2},\ldots,x_{k})\rVert_{Y}:x_{i}\in X,\lVert x_{i}\rVert_{X}\leq 1\}\in[0,\infty].$
We say $\phi$ is continuous if $\lVert\phi\rVert<\infty$.
Definition 2.3.
Let $k\in\mathbb{Z}_{\geq 0}$. We say a map $c\colon X\to Y$ is a (continuous) homogeneous polynomial map of degree $k$ if it is defined by a (continuous) symmetric multilinear map $\overline{c}\colon X^{k}\to Y$ via
$c(x):=\overline{c}(x,x,\ldots,x).$
The map $\overline{c}$ is uniquely determined by $c$ according to the following formula, taken from [Boc70, p. 274], which follows readily from the inclusion–exclusion principle:
$\overline{c}(x_{1},\ldots,x_{k})=\frac{1}{k!}\sum_{\begin{subarray}{c}\epsilon_{1}=0,1\\ \cdots\cdots\\ \epsilon_{k}=0,1\end{subarray}}(-1)^{k-\epsilon_{1}-\ldots-\epsilon_{k}}c\left(\sum_{i=1}^{k}\epsilon_{i}x_{i}\right).$ (1)
Definition 2.4.
A map $f\colon U\to Y$ defined on a neighborhood $U$ of a point $x\in X$ is said to be analytic at $x$ if there exist $r>0$ with $\{z\in X:\lVert z-x\rVert_{X}<r\}\subset U$ and for each $k\in\mathbb{Z}_{\geq 0}$ a continuous homogeneous polynomial map $c_{k}\colon X\to Y$ of degree $k$ such that $\limsup_{k}\lVert\overline{c_{k}}\rVert^{1/k}\leq r^{-1}$ and for all $h\in X$ with $\lVert h\rVert_{X}<r$ one has
$f(x+h)=\sum_{i=0}^{\infty}c_{i}(h).$ (2)
We say $f$ is analytic if $f$ is analytic at each point of its domain $U$.
Many classical results about analytic functions of a single variable extend to the Banach space setting, among which we record for later reference the identity principle.
Theorem 2.5.
Let $U$ be a connected open subset of the Banach space $X$, and let $f,g\colon U\to Y$ be analytic maps. Suppose there exists a nonempty open subset $U_{0}$ of $U$ such that $f(x)=g(x)$ for each $x\in U_{0}$. Then $f$ and $g$ are identically equal on $U$.
Proof.
See [Whi65, Corollary, p. 1080]. ∎
Theorem 2.6.
Let $B$ be a Banach algebra with identity. Then the set $G:=\{b\in B:b\text{ is invertible}\}$ is open and the inverse map $G\to G$, $b\mapsto b^{-1}$ is analytic.
Proof.
This is shown in [Con90, Theorem VII.2.2]. ∎
2.3 Semi-continuous functions
In this section we recall some facts concerning (upper or lower) semicontinuous functions, referring to [Bee93] for proofs.
Definition 2.7.
Let $X$ be a topological space.
  • $(1)$
    A function $g\colon X\to[-\infty,\infty)$ is called upper semicontinuous if for every $\lambda\in\mathbb{R}$ the set $g^{-1}([-\infty,\lambda))$ is open in $X$.
  • $(2)$
    A function $h\colon X\to(-\infty,\infty]$ is called lower semicontinuous if for every $\lambda\in\mathbb{R}$ the set $h^{-1}((\lambda,\infty])$ is open in $X$.
  • $(3)$
    A net $(g_{\alpha})_{\alpha}$ of functions $g_{\alpha}\colon X\to(-\infty,\infty]$ is called almost nondecreasing if for every $\epsilon>0$ there exists an index $\alpha_{0}$ such that for all indices $\alpha\geq\alpha^{\prime}\geq\alpha_{0}$ one has $g_{\alpha}\geq g_{\alpha^{\prime}}-\epsilon$.
  • $(4)$
    A net $(h_{\beta})_{\beta}$ of functions $h_{\beta}\colon X\to[-\infty,\infty)$ is called almost nonincreasing if $(-h_{\beta})_{\beta}$ is an almost nondecreasing net.
Lemma 2.8.
Let $X$ be a topological space.
  1. (1)
    Let $(g_{\alpha})_{\alpha}$ be an almost nonincreasing net which consists of upper semicontinuous functions $g_{\alpha}\colon X\to[-\infty,\infty)$. Then the function $g\colon X\to[-\infty,\infty)$ defined $g(x):=\liminf_{\alpha}g_{\alpha}(x)$ is upper semicontinuous and equal to the pointwise limit of the net $(g_{\alpha})_{\alpha}$.
  2. (2)
    Let $(h_{\beta})_{\beta}$ be an almost nondecreasing net which consists of lower semicontinuous functions $h_{\beta}\colon X\to(-\infty,\infty]$. Then the function $h\colon X\to(-\infty,\infty]$ defined by $h(x):=\limsup_{\beta}h_{\beta}(x)$ is lower semicontinuous and equal to the pointwise limit of the net $(h_{\beta})_{\beta}$.
Proof.
See [Bee93, Theorem 1.3.1]
Lemma 2.9.
Let $X$ be a compact topological space.
  • $(1)$
    Let $g\colon X\to[-\infty,\infty)$ be an upper semicontinuous function. Then $g$ is bounded above and attains its supremum.
  • $(2)$
    Let $h\colon X\to(-\infty,\infty]$ be a lower semicontinuous function. Then $h$ is bounded below and attains its infimum.
Proof.
See [Bee93, Theorem 1.3.4]. ∎
2.4 Order unit spaces
In this section we introduce order unit spaces and state Kadison’s theorem giving a functional representation of complete order unit spaces as Theorem 2.11. We conclude with a discussion of norm completions.
Let $V_{+}$ be a proper cone in a real vector space $V$, i.e. $V_{+}+V_{+}\subset V_{+}$ and $\lambda V_{+}\subset V_{+}$ for all $\lambda\geq 0$, as well as $V_{+}\cap-V_{+}=\{0\}$. A partial order $\leq$ on $V$ is defined by
$x\leq y\iff y-x\in V_{+}.$
We call such a pair $(V,V_{+})$ a partially ordered vector space. It is called archimedean if for every $x\in V$ the set $\{\lambda x:\lambda\geq 0\}$ has an upper bound only if $x\leq 0$. An element $v\in V_{+}$ is called an order unit if for every $x\in V$ there exists $\lambda>0$ such that $x\leq\lambda v$. A triple $(V,V_{+},v)$ is called an order unit space if $(V,V_{+})$ is an archimedean partially ordered vector space with order unit $v$.
Let $(V,V_{+},v)$ be an order unit space. Its order unit norm is the norm $\lVert\cdot\rVert_{v}$ on $V$ defined by
$\lVert x\rVert_{v}:=\{\inf\lambda>0:-\lambda v\leq x\leq\lambda v\}.$
Its open cone is the $\lVert\cdot\rVert_{v}$-interior $V^{\circ}_{+}$ of $V_{+}$, given by
$V^{\circ}_{+}=\{x\in V:\epsilon v\leq x\text{ for some }\epsilon>0\}.$
The elements of $V^{\circ}_{+}$ are precisely the order units of the partially ordered vector space $(V,V_{+})$. If $v^{\prime}\in V^{\circ}_{+}$, then the norm $\lVert\cdot\rVert_{v^{\prime}}$ equivalent to $\lVert\cdot\rVert_{v}$. We call $(V,V_{+},v)$ a complete order unit space if $(V,\lVert\cdot\rVert_{v})$ is a Banach space.
Example 2.10.
Let $K$ be a compact convex subset of a locally convex space $E$. We write $A(K)$ for the real linear space of continuous affine functions $a\colon K\to\mathbb{R}$. We give $A(K)$ the pointwise ordering, so that $A(K)_{+}=\{a\colon K\to[0,\infty)\text{ affine continuous}\}$. The triple $(A(K),A(K)_{+};\mathbbm{1}_{K})$ is a complete order unit space, where $\mathbbm{1}_{K}:K\to\mathbb{R}$ is identically equal to $1$.
We now recall a fundamental theorem of Kadison, stating that each complete order unit space is isomorphic to $A(K)$ for a certain compact convex subset $K$ of $V^{*}$, called the state space.
Let $(V,V_{+},v)$ be an order unit space. Its state space is the set $K$ of positive linear functionals $\rho\colon V\to\mathbb{R}$ such that $\rho(v)=1$. Since each state is $\lVert\cdot\rVert_{v}$-continuous, $K$ is a subset of the continuous dual $V^{*}$ of $(V,\lVert\cdot\rVert_{v})$. We endow $V^{*}$ with the wk*-topology. It follows from the Banach–Alaoglu theorem that $K$ is a wk*-compact convex subset of $V^{*}$.
Theorem 2.11 (Kadison).
Let $(V,V_{+},v)$ be a complete order unit space. For $x\in V$ define $\tilde{x}\colon K\to\mathbb{R}$ by $\tilde{x}(\rho):=\rho(x)$ for all $\rho\in K$. Then there exists an isomorphism of order unit spaces
$(V,V_{+},v)$ $\stackrel{{\scriptstyle\sim}}{{\to}}(A(K),A(K)_{+},\mathbbm{1}_{K}),$ (3)
$x$ $\mapsto\tilde{x}.$
Proof.
We refer to [Alf71, Theorem II.1.8]. ∎
Remark 2.12.
If $V$ is a, possibly non-complete, order unit space, then the assignment $x\mapsto\tilde{x}$ defines an isometric isomorphism of $V$ onto a norm dense subspace of $A(K)$, namely the space $A(K,V^{*})$ consisting of all functions on $K$ that extend to a continuous real-valued affine function on $V^{*}$ (see again [Alf71, Theorem II.1.8]). In particular, the norm completion $\hat{V}$ of $V\cong A(K,V^{*})$ is identified as $\hat{V}\cong A(K)$.
Of fundamental importance in the theory of compact convex sets is the Krein–Milman theorem, ensuring the existence of extremal points. Recall that if $L$ is a convex subset of a vector space $E$, then an extremal point of $L$ is a point $\rho\in L$ such that for all $10<\lambda<$ 1 and $\sigma,\tau\in L$ with $\rho=(1-\lambda)\sigma+\lambda\tau$ one has $\sigma=\tau=\rho$, i.e., $\rho$ is not a proper convex combination of points in $L$ different from $\rho$. The set of extremal points of $L$ is denoted $\delta_{e}L$.
Theorem 2.13 (Krein–Milman).
Let $L$ be a compact convex subset of a locally convex space. Then $L$ is the closed convex hull of the set $\delta_{e}L$ of extremal points of $L$.
Proof.
We refer to [Con90, Theorem V.7.4]. ∎
An extremal point of the state space $K$ of an order unit space $(V,V_{+},v)$ is called a pure state. So, the set of pure states of $V$ is denoted $\partial_{e}K$. It follows from the Krein–Milman theorem that if $a,a^{\prime}\in A(K)$ agree on $\delta_{e}K$, then $a=a^{\prime}$ holds; see Lemma 3.8 for a more general statement.
We conclude with a lemma on the completion of an order unit space, which will be applied in Section 5.8.
Lemma 2.14.
Let $(V,V_{+},v)$ be an order unit space. Let $\hat{V}$ be the completion of the normed space $(V,\lVert\cdot\rVert_{v})$, and let $\hat{V}_{+}$ be the closure of $V_{+}$ in $\hat{V}$. Then $(\hat{V},\hat{V}_{+},v)$ is a complete order unit space and the order unit norm on $\hat{V}$ coincides with the extension of the norm $\lVert\cdot\rVert_{v}$ on $V$ to its completion.
Proof.
By Remark 2.12, we may assume that $V=A(K,E)$ for some locally convex space $E$ and compact convex subset $K$ of $E$, so that $\hat{V}=A(K)$. Let us verify that the closure of $A(K,E)_{+}$ inside $A(K)$ equals $A(K)_{+}$. Plainly the closure of $A(K,E)_{+}$ is contained in $A(K)_{+}$. Conversely, let $x\in A(K)_{+}$. Since $A(K,E)$ is norm dense in $A(K)$ we can find a sequence $(x_{n})_{n}$ in $A(K,E)$ converging to $x$ in $\lVert\cdot\rVert_{v}$, i.e. uniformly on $K$. Therefore, the sequence $(\rho_{n})_{n}$ defined by $\rho_{n}:=\inf_{K}x_{n}$ converges to $\inf_{K}x\geq 0$. It follows that the sequence $(y_{n})_{n}$ defined by $y_{n}:=x_{n}+\max(0,-\inf x_{n})v$ belongs to $A(K,E)_{+}$ and converges to $x+\max(0,-\inf_{K}x)v=x$. We conclude that $x$ is contained in the closure of $A(K,E)_{+}$, as desired.
The final assertion about coincidence of the two norms on $\hat{V}_{+}$ holds because both norms agree with the supremum norm on $A(K)$. ∎
In the forthcoming section, we will identify $x$ with $\tilde{x}$ and write $x(\rho)$ instead of $\tilde{x}(\rho)=\rho(x)$, for $x\in V$ and $\rho\in K$.
2.5 Gauges and Thompson’s metric
Let $(V,V_{+},v)$ be an order unit space. On the open cone $C=V^{\circ}_{+}$ there exists a natural metric $d_{T}$ called Thompson’s metric. We will show that $d_{T}$ is locally bi-Lipschitz to the metric associated with the order unit norm $\lVert\cdot\rVert_{v}$. Moreover, we define gauge functions, which measure or gauge how small or large one element of $C$ is compared to another.
In the following definition, observe that since $C$ consists of the order units of $(V,V_{+})$, for any two $x,y\in C$ there exists $\lambda>0$ such that $x\leq\lambda y$.
Definition 2.15.
Let $(V,V_{+},v)$ be an order unit space and $C=V_{+}^{\circ}$ its open cone.
  1. (1)
    We define Thompson’s metric $d_{T}\colon C\times C\to\mathbb{R}_{\geq 0}$ by
    $d_{T}(x,y)=\inf\{\lambda>0:x\leq\lambda y\text{ and }y\leq\lambda x\}.$ (4)
  2. (2)
    We define the gauge-functions $L\colon C\times C\to\mathbb{R}_{>0}$ and $M\colon C\times C\to\mathbb{R}_{>0}$ by
    $m(x,y):=\sup\{\lambda>0:\lambda y\leq x\}\in\mathbb{R}_{>0},$ (5)
    and
    $M(x,y):=\inf\{\mu>0:x\leq\mu y\}\in\mathbb{R}_{>0}.$ (6)
In terms of the gauge functions, Thompson’s metric is given by
$d_{T}(x,y)=\log\max\{M(x,y),M(y,x)\}.$ (7)
One deduces that $d_{T}$ satisfies the triangle inequality and is nonnegative from the inequality
$M(x,z)\leq M(x,y)M(y,z),$ (8)
valid for all $x,y,z\in C$.
Example 2.16.
Using Kadison’s Theorem 2.11, we may view elements $x,y\in C$ as continuous functions from the state space $K$ to $\mathbb{R}_{>0}$. Their quotient $x/y\colon K\to\mathbb{R}_{>0}$, $\psi\mapsto x(\psi)/y(\psi)$ is a continuous function on $K$ as well. The minimum and maximum values of $x/y$ over $K$ are given by the gauge function values $m(x,y)$ resp. $M(x,y)$, that is
$m(x,y)=\min_{\psi\in K}\frac{x(\psi)}{y(\psi)}\text{ and }M(x,y)=\max_{\psi\in K}\frac{x(\psi)}{y(\psi)}.$ (9)
The gauge functions have the following elementary properties.
Lemma 2.17.
Let $x,y\in C$ and $\lambda>0$. Then
$m(y,x)$ $=M(x,y)^{-1},$ (10)
$M(\lambda x,y)$ $=\lambda M(x,y)$ (11)
$M(x,\lambda y)$ $=\lambda^{-1}M(x,y)$ (12)
$m(y,\lambda x)$ $=\lambda^{-1}m(y,x)$ (13)
$m(\lambda y,x)$ $=\lambda m(y,x)$ (14)
$d_{T}(\lambda x,\lambda y)$ $=d_{T}(x,y).$ (15)
Proof.
Left to the reader. ∎
Lemma 2.18.
Let $\lambda>0$. For all $x,y\in C$ with $x,y\geq\lambda^{-1}v$ we have
$d_{T}(x,y)\leq\lambda\lVert x-y\rVert_{v}.$
Proof.
For each state $\psi$ on $(V,v)$ we have, because $y(\psi)\geq m(y,v)\geq\lambda^{-1}$, that
$\frac{x(\psi)}{y(\psi)}=1+\frac{x(\psi)-y(\psi)}{y(\psi)}\leq 1+\frac{\lVert x-y\rVert_{v}}{m(y,v)}\leq 1+\lambda\lVert x-y\rVert_{v}.$
This shows that $M(x,y)\leq 1+\lambda\lVert x-y\rVert_{v},$ and the same holds with $x$ and $y$ interchanged. We conclude that
$d_{T}(x,y)\leq e^{d_{T}(x,y)}-1=\max\{M(x,y),M(y,x)\}-1\leq\lambda\lVert x-y\rVert_{v}.$
Lemma 2.19.
Let $\lambda>0$. For all $x,y\in C$ with $x,y\leq\lambda v$ we have
$\lVert x-y\rVert_{v}\leq\lambda d_{T}(x,y).$
Proof.
Let $\psi$ be a state on $(V,v)$. We have $y(\psi)\geq e^{-d_{T}(x,y)}x(\psi)$ because $e^{d_{T}(x,y)}y\geq M(x,y)y\geq x$. Using that $1-e^{-a}\leq a$ for each $a\in\mathbb{R}$, we obtain
$x(\psi)-y(\psi)\leq(1-e^{-d_{T}(x,y)})x(\psi)\leq\lambda d_{T}(x,y).$
By interchanging $x$ and $y$ we find that $|x(\psi)-y(\psi)|\leq\lambda d_{T}(x,y)$. The desired inequality follows since $\psi$ is an arbitrary state on $(V,v)$. ∎
As a consequence of the two preceding lemmata, Thompson’s metric $d_{T}$ and the metric associated with the order unit norm $\lVert\cdot\rVert_{v}$ are bi-Lipschitz with constant $\lambda$ on the Thompson’s metric ball $\{x\in V:d_{T}(v,x)\leq e^{\lambda}\}=[\lambda^{-1}v,\lambda v]$ around $v$ of radius $e^{\lambda}$ for each $\lambda>0$, hence these metrics are locally bi-Lipschitz. In particular, $d_{T}$ and $\lVert\cdot\rVert_{v}$ induce the same topology on $C$.
2.6 Gauge-reversing bijections
In this section we define gauge-preserving bijections and gauge-reversing bijections. We show they are isometries for Thompson’s metric and, consequently, are locally bi-Lipschitz with respect to the order unit norms. Moreover, we introduce $d_{T}$-symmetries, and state that these are gauge-reversing bijections.
Definition 2.20.
Let $(X,\leq)$ and $(Y,\leq)$ be partially ordered sets, and let $\Phi\colon X\to Y$ be a map. We say that $\Phi$ is order-preserving (resp. order-reversing) if for all $x_{1},x_{2}\in X$ with $x_{1}\leq x_{2}$ one has $\Phi(x_{1})\leq\Phi(x_{2})$ (resp. $\Phi(x_{1})\geq\Phi(x_{2})$). We say $\Phi$ is an order-isomorphism (resp. order-anti-isomorphism if $\Phi$ is bijective and both $\Phi$ and $\Phi^{-1}$ are order-preserving (resp. order-reversing).
In the remainder of this section, $(V,V_{+},v)$ and $(W,W_{+},w)$ denote order unit spaces, and $C=V_{+}^{\circ}$ and $D=W_{+}^{\circ}$ their open cones.
Definition 2.21.
Let $\Phi\colon C\to D$ be a bijection. We say that $\Phi$ is homogeneous of degree $1$ (resp. homogeneous of degree $-1$) if for all $x\in C$ and $\lambda\in(0,\infty)$ one has $\Phi(\lambda x)=\lambda\Phi(x)$ (resp. $\Phi(\lambda x)=\lambda^{-1}\Phi(x)$). We call $\Phi$ gauge-preserving if for all $x,y\in C$ one has
$m(\Phi(x),\Phi(y)=m(y,x)$ (16)
and gauge-reversing if all $x,y\in C$ satisfy
$m(\Phi(x),\Phi(y))=m(x,y).$ (17)
Since the functions $M$ and $m$ are multiplicatively inverse by (31), one arrives at an equivalent definition if condition (16) is replaced by $M(\Phi(x),\Phi(y))=M(x,y)$ and condition (17) by $M(\Phi(x),\Phi(y))=M(y,x)$.
Proposition 2.22.
For a bijection $\Phi\colon C\to D$, the following statements are equivalent:
  1. (1)
    $\Phi$ is gauge-reversing;
  2. (2)
    $\Phi$ is homogeneous of degree $-1$ and an order-anti-isomorphism.
Proof.
See [LRW25, Lemma 2.5]. ∎
Proposition 2.23.
For a bijection $\Psi\colon C\to D$, the following statements are equivalent:
  1. $(1)$
    $\Psi$ is gauge-preserving;
  2. $(2)$
    $\Psi$ is homogeneous of degree $1$ and an order-isomorphism.
Proof.
The proof is similar to Proposition 2.22 and therefore omitted. ∎
Theorem 2.24.
Let $\Psi\colon C\to D$ be a gauge-preserving bijection sending $v$ to $w$. Then $\Psi$ extends to an isomorphism $V\to W$ of the order unit spaces $(V,V_{+},v)$ and $(W,W_{+},w)$.
Proof.
This is proved as Theorem B in [NS77] and [Sch78]. ∎
We will only have need for the previous theorem in the special case that $(V,V_{+},v)$ of $(W,W_{+},w)$ are complete and there exists a gauge-reversing bijection $\Phi\colon C\to D$. In order to make our exposition more self-contained, we will provide in Section 4.2 an independent proof of Theorem 2.24 under an additional assumption (namely that the cone $C_{usc}$ be strongly atomic, cf. Section 3.3), which is satisfied in our special case by Theorem 3.20.
Proposition 2.25.
Each gauge-reversing or gauge-preserving bijection $\Phi\colon C\to D$ is a $d_{T}$-isometry.
Proof.
Suppose first that $\Phi$ is gauge-reversing. Then for all $x,y\in C$ we have that
$d_{T}(\Phi(x),\Phi(y))$ $=\log\max\{M(\Phi(x),\Phi(y)),M(\Phi(y),\Phi(x))\}$
$=\log\max\{M(y,x),M(x,y)\}$
$=d_{T}(x,y),$
hence $\Phi$ is a $d_{T}$-isometry. The proof for a gauge-preserving bijection is similar. ∎
Definition 2.26.
We define a $d_{T}$-symmetry $S_{x}\colon C\to C$ at a point $x\in C$ to be an involutive $d_{T}$-isometry having $x$ as an isolated fixed point.
Theorem 2.27.
Let $S_{x}:C\to C$ be a $d_{T}$-symmetry at some point $x\in C$. Then $S_{x}$ is a gauge-reversing bijection.
Proof.
This is shown in [LRW25, Theorem 3.5]. ∎
Proposition 2.28.
Let $\lambda>0$. For all $x,y\in C$ with $x,y\geq\lambda^{-1}v$ one has
$\lVert\Phi(x)-\Phi(y)\rVert_{w}\leq\lambda^{2}\lVert x-y\rVert_{v}.$ (18)
Proof.
In Lemma 2.18 we showed that $d_{T}(x,y)\leq\lambda\lVert x-y\rVert_{v}$. Likewise, since $\Phi(x),\Phi(y)\geq\lambda w$ we have by Lemma 2.19 that $\lVert\Phi(x)-\Phi(y)\rVert_{w}\leq\lambda d_{T}(\Phi(x),\Phi(y))$. Because $\Phi$ is a $d_{T}$-isometry by Proposition 2.25, we conclude that
$\lVert\Phi(x)-\Phi(y)\rVert_{w}\leq\lambda d_{T}(\Phi(x),\Phi(y))=\lambda d_{T}(x,y)\leq\lambda^{2}\lVert x-y\rVert_{v}.$
Theorem 2.29.
The map $\Phi\colon(C,\lVert\cdot\rVert_{v})\to(D,\lVert\cdot\rVert_{w})$ is bi-Lipschitz on Thompson’s metric balls around $v$ and $w$. In particular, $\Phi$ is locally bi-Lipschitz.
Proof.
Fox fixed $\lambda>0$, the maps $\Phi$ and $\Phi^{-1}$ restrict to bijections between the Thompson’s metric balls $[\lambda^{-1}v,\lambda v]$ and $[\lambda^{-1}w,\lambda w]$, and these restrictions are Lipschitz continuous with constant $\lambda^{2}$ by Proposition 2.28. Since the norm interiors of $[\lambda^{-1}v,\lambda v]$ for $\lambda>0$ cover $C$, the corollary follows. ∎
2.7 Jordan algebras
In this section we first define classical, i.e. linear, Jordan algebras. We then provide an alternative axiomatic framework devised by McCrimmon, and show each quadratic Jordan algebra is a linear Jordan algebra. The theory in this section is purely algebraic and is valid over any field of characteristic $\neq 2$ with at least four elements. However, for simplicity we work over the field of real numbers $\mathbb{R}$. Throughout this paper, all our (Jordan) algebras are assumed to have an identity element.
Definition 2.30.
A Jordan algebra is a real linear space $A$ equipped with a commutative, but not necessarily associative, $\mathbb{R}$-bilinear product that satisfies the so-called Jordan-identity
$x\circ(y\circ x^{2})$ $=(x\circ y)\circ x^{2}$ (J)
and has a unit element $e$, i.e. $x\circ e=x$, for all $x,y\in A$.
Example 2.31.
Let $(A,\cdot)$ be an associative real algebra with identity $e$. Define on $A$ the following Jordan product $\circ$ by
$x\circ y:=\tfrac{1}{2}(x\cdot y+y\cdot x).$
Then $(A,\circ)$ is a Jordan algebra with identity element $e$.
For a real vector space $V$, we denote $L(V)$ the real algebra of $\mathbb{R}$-linear endomorphisms of $V$.
Definition 2.32.
Let $(A,\circ)$ be a Jordan algebra, and let $x\in A$ be an element.
  • $(1)$
    We define the linear representation of $x$ to be the endormorphism $T(x)\in L(A)$ given by
    $T(x)(y):=x\circ y=y\circ x.$
  • $(2)$
    The quadratic representation of $x$ is defined in terms of the linear representation as
    $U(x):=2T(x)^{2}-T(x^{2})\in L(A).$ (19)
The map $T\colon A\to L(A)$ is linear, but the quadratic representation map $U\colon A\to L(A)$ is a quadratic map in the sense of the following definition.
Definition 2.33.
Let $A$ and $B$ be real vector spaces. We say a map $U\colon A\to B$ is quadratic if it is homogeneous of degree $2$, i.e. $U(\lambda x)=\lambda^{2}U(x)$ for all $\lambda\in\mathbb{R}$ and $x\in A$, and the map $U:A\times T\to B$ defined by
$U(x,y):=\tfrac{1}{2}(U(x+y)-U(x)-U(y))\in B$
is bilinear.
Since $U(x,e)y=x\circ y$ for all $x,y\in A$, the bilinear product $x\circ y$ can be reconstructed from the quadratic product $U(x)y=2x\circ(x\circ y)-x^{2}\circ y$. In fact, McCrimmon has expressed the axioms for a Jordan algebra in terms of the quadratic representation.
Definition 2.34.
A (unital) quadratic Jordan algebra is a triple $(A,U,e)$ consisting of a real vector space $A$, a vector $e\in A$ and a quadratic map $U\colon A\to\operatorname{End}(A)$ satisfying the following axioms. Let $U(\cdot,\cdot)\colon A\to\operatorname{End}(A)$ be the symmetric bilinear map given by $U(x,y)=\frac{1}{2}(U(x+y)-U(x)-U(y))$. Then for all $x,y,z\in A$, we impose that
$U(e)$ $=\operatorname{Id}_{A},$ (QJ1)
$U(x)U(y,z)x$ $=U(U(x)y,x)z,$ (QJ2)
$U(U(x)y)$ $=U(x)U(y)U(x).$ (QJ3)
If $A$ is a classical Jordan algebra with unit element $e$ and quadratic map $U:A\to L(A)$, then $(A,U,e)$ is a quadratic Jordan algebra, as is shown in [Jac69, §1.2]. Conversely, a quadratic Jordan algebra $(A,U,e)$ gives rise to a classical Jordan algebra with identity element $e$ and quadratic representation map $U$. This result is proved in [Jac69, §1.3 and §1.4], among a list of other identities. To make our exposition more self-contained, we reproduce his derivation in the following theorem.
Theorem 2.35.
Let $(A,U,e)$ be a quadratic Jordan algebra. Then $A$ is a Jordan algebra with unit element $e$ in the product $a\circ b:=U(a,b)e$.
Proof.
It is clear that the product $\circ$ is bilinear and commutative. For every $a\in A$ set
$T(a)=U(a,e)\in\operatorname{End}(A).$
Take $x=e$ in (QJ2) to find using (QJ1) that
$T(y)z=U(y,e)z=U(y,z)e=y\circ z,$ (20)
so $T(y)$ is the operation of multiplication by $y$. In particular, $T(e)=U(e,e)=U(e)=\operatorname{Id}_{A}$, so $e$ is the multiplicative identity element. Moreover, $U(x)e=U(x,x)e=x\circ x=:x^{2}$.
Next we show that for every $a\in A$ we have
$U(a)=2T(a)^{2}-T(a^{2}).$
Let $\lambda$ be real number, and take $x=a+\lambda e$ and $y=e$ in (QJ3). This gives, since $U(e)=\operatorname{id}_{A}$, that
$U(a+\lambda e)^{2}=U((a+\lambda e)^{2}),$
which expands, using (QJ1) again, as
$(U(a)+2\lambda U(a,e)+\lambda^{2}\operatorname{Id}_{A})^{2}=U(a^{2}+2\lambda a+\lambda^{2}e).$
Comparison of coefficients at $\lambda^{2}$ gives
$2U(a)+4U(a,e)^{2}=4U(a)+2U(a^{2},e).$
Dividing by $2$ and substitute $T(a)=U(a,e)$ and $T(a^{2})=U(a^{2},e)$ to obtain $U(a)=2T(a)^{2}-T(a^{2})$.
Finally, we prove the Jordan identity by showing that $T(x)$ and $T(x^{2})$ commute. Take $z=e$ in (QJ2) to find using (20) that
$U(x)U(x,e)y=U(x)U(y,e)x=U(U(x)y,x)e=U(x,e)U(x)y,$
i.e. $U(x)$ and $U(x,e)=T(x)$ commute. Since we have shown that $T(x^{2})=2T(x)^{2}-U(x)$, it follows that $T(x)$ and $T(x^{2})$ commute. ∎
If $A$ and $B$ are Jordan algebras, then a linear map $\phi:A\to B$ is a Jordan homomorphism if and only if $\phi(1_{A})=1_{B}$ and $U_{B}(\phi(x))\circ\phi=\phi\circ U_{A}(x)$. It follows that the categories of Jordan algebras and quadratic Jordan algebras are equivalent (see also [Jac69, §1.4]).
Definition 2.36.
An element $x\in A$ is called invertible if its quadratic representation $U(x)$ is invertible in the Banach algebra $L(A)$.
Remark 2.37.
This notion of invertibility agrees with the one commonly used in the theory of linear Jordan algebras, where an element $x\in A$ is called invertible if there exists $y\in A$ such that $x\circ y=e$ and $x^{2}\circ y=x$. Indeed, such element $y$ exists if and only if $U(x)$ is invertible, in which case $y=U(x)^{-1}$. The element $x^{-1}:=U(x)^{-1}x$ is called the inverse of $x$, and has quadratic representation $U(x^{-1})=U(x)^{-1}$.
2.8 JB-algebras
In this section we define JB-algebras, which are Jordan algebras with a compatible norm. The compatibility conditions are customarily stated in terms of the linear product, however, may equivalently be formulated in terms of the quadratic representation operators. This observation allows to present a definition of JB-algebra referencing only the quadratic product.
Definition 2.38.
[…]
$\lVert x\circ y\rVert$ $\leq\lVert x\rVert\lVert y\rVert,$ (21)
$\lVert x^{2}\rVert$ $=\lVert x\rVert^{2},$ (22)
$\lVert x^{2}\rVert$ $\leq\lVert x^{2}+y^{2}\rVert.$ (23)
It follows from (22) that $\lVert e\rVert=1$ provided $A\neq 0$. Each JB-algebra is an order unit space per the following proposition.
Proposition 2.39.
[…]
  • $(1)$
[…]
  • $(2)$
    For every $x\in A$ the map $U(x):A\to A$ is positive and has norm $\lVert U(x)\rVert=\lVert x\rVert^{2}_{e}$.
Proof.
For (1) see [HOS84, Lemma 3.3.7 and 3.3.10]; for (2) see [AS03, Thm. 1.25]. ∎
The following theorem provides a converse.
Theorem 2.40.
Let $(A,A_{+},e)$ be a complete order unit space which is simultaneously a Jordan algebra that has $e$ as its identity element. Suppose for each $x\in A$, the quadratic representation $U(x)\in L(A)$ of $x$ is positive and has $\lVert\cdot\rVert_{e}$-operator norm $\lVert U(x)\rVert_{e}=\lVert x\rVert^{2}$. Then $A$ is a JB-algebra in the order unit norm $\lVert\cdot\rVert_{e}$ with cone of squares $A_{+}=\{x\in A:x^{2}\}$.
Proof.
From [AS01, Lemma 1.80] it follows that $\lVert x\circ y\rVert_{e}\leq\lVert x\rVert_{e}\lVert y\rVert_{e}$ for all $x,y\in A$ and that all squares are positive. Conversely, by [ibidem, Lemma 1.73] each positive element is a square. Since $\lVert\cdot\rVert_{e}$ is monotone on $A_{+}=\{x^{2}:x\in A\}$, we have $\lVert x^{2}\rVert_{e}\leq\lVert x^{2}+y^{2}\rVert_{e}$ for all $x,y\in V$. Finally, according to [ibidem, Theorem 1.81] we have $\lVert x^{2}\rVert_{e}=\lVert x\rVert_{e}^{2}$ for each $x\in A$. We have shown that the norm $\lVert\cdot\rVert_{e}$ satisfies (21), (22) and (23). We conclude that $A$ is a JB-algebra in the order unit norm $\lVert\cdot\rVert_{e}$ with cone of squares $A_{+}$. ∎
Theorem 2.35, Proposition 2.39 and Theorem 2.40 combine to show the equivalence of the standard definition of JB-algebras encountered in the literature, to the following alternative definition, which we will employ exclusively in the remainder of this article.
Definition 2.41.
A JB-algebra is a complete order unit space $(A,A_{+},e)$ together with a quadratic map $U\colon A\to L(A)$ satisfying (QJ1-QJ3) such that for every $x\in A$ the quadratic representation $U(x)\in L(A)$ is positive and has $\lVert\cdot\rVert_{e}$-operator norm $\lVert U(x)\rVert_{e}=\lVert x\rVert^{2}_{e}$.
The only if direction of Theorem 1.1 is a known result in the theory of JB-algebras.
Theorem 2.42.
Let $(A,A_{+},e,U)$ be a quadratic JB-algebra, and let $\Omega:=A_{+}^{\circ}$ be its open cone. For $x\in\Omega$ denote $i(x):=x^{-1}(=U(x)^{-1}x)$ the inverse of $x$. Then the map $i\colon\Omega\to\Omega$ is Fréchet-differentiable with derivative
$Di(x)=-U(x)^{-1}.$
Moreover, $i$ is an involutive gauge-reversing bijection satisfying $i(e)=e$ and $Di(e)=-\operatorname{Id}_{A}$.
Proof.
The Fréchet-derivative of $i$ is determined in [FK94, Prop. II.3.3(i)] for Euclidean Jordan algebras; the proof is also valid for JB-algebras.
It is clear that $i$ is involutive and homogeneous of degree $-1$. In [AS03, Lemma 1.31] it is shown that $i$ is order-reversing. Since $i=i^{-1}$ this entails that $i$ is an order-anti-isomorphism. We conclude that $i$ is an involutive gauge-reversing bijection. Finally, $U(e)=\operatorname{Id}_{A}$ yields that $i(e)=U(e)^{-1}e=e$ and $Di(e)=-U(e)^{-1}=-\operatorname{Id}_{A}$. ∎
3 Extended cones of semicontinuous affine functions
It will be useful to extend the given gauge-reversing bijection to certain larger cones of semicontinuous affine functions. We will introduce these cones in Section 3.1, and construct the extension in Section 3.2. The technical advantage of these extended cones is that they possess a wealth of extremal structure, as will be discussed in Section 3.3.
3.1 Extension of the cones
Throughout this section, we let $(V,V_{+},v)$ be a complete order unit space. We let $K$ be its state space, always considered in the wk*-topology. Recall Kadison’s Theorem 2.11 stating that there exists an isomorphism $(V,V_{+},v)\cong(A(K),A(K)_{+},\mathbbm{1}_{K})$ sending $x\in V$ to the function $\tilde{x}:K\to\mathbb{R}$ given by $\tilde{x}(\rho):=\rho(x)$. We will identify each element $x\in V$ with the affine continuous function $\tilde{x}$ on $K$ it defines, thus dropping the tilde in $\tilde{x}$. Hence, we will write $x(\rho)$ for the value $\rho(x)=\tilde{x}(\rho)$.
We denote the open cone by $C:=V_{+}^{\circ}$. Since $C=\{x\in V:x\geq\delta v\text{ for some }\delta>0\}$ and $K$ is compact, we have
$C=\{a\colon K\to(0,\infty)\text{ affine continuous}\}.$
The set of all $[0,\infty]$-valued affine functions on $K$ is denoted
$A(K)_{+}^{\infty}:=\{g\colon K\to[0,\infty]\text{ affine}\}.$
We endow $A(K)_{+}^{\infty}$ with the pointwise ordering given by $a\leq b$ if $a(\rho)\leq b(\rho)$ for every $\rho\in K$. We remark that $A(K)_{+}^{\infty}$ is a monotone complete partially ordered set: each nonincreasing net in $A(K)^{\infty}_{+}$ has an infimum, and each nondecreasing net in $A(K)_{+}^{\infty}$ has a supremum.
Let us also define a transitive relation $\prec$ on $A(K)^{\infty}_{+}$ by $a\prec b$ if $a(\rho)<b(\rho)$ for every $\rho\in K$. Note that for $a,b\in A(K)_{+}$ we have $a\prec b$ if and only if $b-a\in C$, whereas $a\leq b$ if and only if $b-a\in A(K)_{+}$.
Moreover, on $A(K)_{+}^{\infty}$ we define a pointwise addition and multiplication by scalars in $[0,\infty]$, using the convention that $0\cdot\infty=0$. In this way, $A(K)_{+}^{\infty}$ becomes a semimodule over the semiring $[0,\infty]$.
We now introduce two cones in $A(K)_{+}^{\infty}$ containing $C$, which will figure prominently in the sequel:
$C_{usc}:=\{g\colon K\to[0,\infty)\mid\text{ affine upper semicontinuous}\}$ (24)
and
$C_{lsc}:=\{h\colon K\to(0,\infty]\mid\text{ affine lower semicontinuous}\}.$ (25)
We remark that $C_{lsc}$ is closed under addition and multiplication by scalars in $(0,\infty]$ and that $C_{usc}$ is closed under addition and multiplication by scalars in $[0,\infty)$. Furthermore, we have that
$C_{usc}\cap C_{lsc}=\{g\colon K\to(0,\infty)\text{ affine continuous}\}=C.$
Example 3.1.
Let $F$ be a nonempty closed face of $K$. Define the function $I^{\infty}_{F}\colon K\to[0,\infty]$ by
$I^{\infty}_{F}(\rho)=\begin{cases}1&\text{if }\rho\in F,\\ \infty&\text{if }\rho\not\in F.\end{cases}$ (26)
We have $I^{\infty}_{F}\in C_{usc}$ with $\min I^{\infty}_{F}=1$. In particular, if $\psi\in\delta_{e}K$ is a pure state, then $I^{\infty}_{\{\psi\}}\in C_{usc}$.
Remark 3.2.
It would have been equivalent to define $C_{usc}$ and $C_{lsc}$ as cones of affine semicontinuous functions on the dual cone of $V_{+}$. Indeed, let $V^{*}_{+}$ be the dual cone of $V_{+}$ in $V^{*}$, consisting of all positive linear functions on $V$, so that $K=\{\rho\in V^{*}_{+}:\rho(v)=1\}$. Restriction defines an order- and $[0,\infty]$-semimodule-isomorphism
$\{\tilde{g}\colon V^{*}_{+}\to[0,\infty]\text{ affine}\}$ $\stackrel{{\scriptstyle\sim}}{{\to}}A(K)^{\infty}_{+},$ (27)
$\tilde{g}$ $\mapsto\tilde{g}|_{K},$ (28)
because each $g\in A(K)^{\infty}_{+}$ has a unique affine extension $\tilde{g}\colon V^{*}_{+}\to[0,\infty]$ given by
$\tilde{g}(\rho)=\begin{cases}0&\text{if }\rho=0,\\ \rho(v)g(\rho(v)^{-1}\rho)&\text{if }\rho\neq 0.\end{cases}$
Moreover, $g$ is upper or lower semicontinuous if and only if $\tilde{g}$ is so, whence the isomorphism (27) given by $\tilde{g}\mapsto\tilde{g}|_{K}$ restricts to order-isomorphisms
$\{\tilde{g}:V^{*}_{+}\to[0,\infty)\text{ affine upper semicontinuous}\}\stackrel{{\scriptstyle\sim}}{{\to}}C_{usc}$
and
$\{\tilde{g}:V^{*}_{+}\to(0,\infty]\text{ affine lower semicontinuous}\}\stackrel{{\scriptstyle\sim}}{{\to}}C_{lsc}.$
It follows from this alternative description of $C_{usc}$ and $C_{lsc}$, that if $v^{\prime}\in C$ is a second order unit and $C^{\prime}_{lsc}$ and $C^{\prime}_{usc}$ are the cones defined by (25) and (24) but with $K$ replaced by the state space $K^{\prime}:=\{\rho\in V^{*}_{+}:\rho(v^{\prime})=1\}$ of $(V,V_{+},v^{\prime})$, then there exist canonical cone isomorphisms $C_{usc}\to C^{\prime}_{usc}$ and $C_{lsc}\to C^{\prime}_{lsc}$ given by $g\mapsto\tilde{g}|_{K^{\prime}}$ which respect all structure. In conclusion, the cones $C_{usc}$ and $C_{lsc}$ are independent of the choice of the order unit $v\in C$ for $(V,V_{+})$.
Our next theorem shows that $C_{lsc}$ is the closure of $C$ inside $A(K)_{+}^{\infty}$ under the formation of suprema of nondecreasing nets. Similarly, $C_{usc}$ consists precisely of the infima of nonincreasing nets in $C$. We will deduce these approximation results from two lemmata.
Lemma 3.3.
Let $K$ be the state space of a complete order unit space $(V,V_{+},v)$. Consider an affine upper semicontinuous function $g\colon K\to[-\infty,\infty)$, and set
$\mathfrak{A}:=\{a\in A(K):a\succ g\}.$
  1. (1)
    For all $a_{1},a_{2}\in\mathfrak{A}$ there exists $a\in\mathfrak{A}$ with $a\leq a_{i}$ for $i\in\{1,2\}$.
  2. (2)
    For every $\rho\in K$ we have $g(\rho)=\inf\{a(\rho):a\in\mathfrak{A}\}$.
Proof.
Part (1) follows from [Alf71, Cor. I.1.4] and part (2) follows from [ibidem, Cor. I.1.2]. ∎
We now make a remark about our language concerning nets. Let $E$ be a nonempty subset of $A(K)^{\infty}_{+}$ such that for any two $h_{1},h_{2}\in E$ there exists $h\in E$ with $h\geq h_{i}$ for $i\in\{1,2\}$. Then the inclusion map $E\to A(K)^{\infty}_{+}$ is a nondecreasing net indexed by the directed partially ordered set $(E,\leq)$. By abuse of language, we will speak of the nondecreasing net $E$. Similarly, if $D$ is a nonempty subset of $A(K)^{\infty}_{+}$ such that for any two $g_{1},g_{2}\in D$ there exists $g\in D$ with $g\leq g_{i}$ for $i\in\{1,2\}$, then we will view $D$ as nonincreasing net indexed by itself.
Lemma 3.4.
Let $(V,V_{+},v)$ be a complete order unit space, with open cone $C:=V_{+}^{\circ}$.
  1. (1)
    Each nonincreasing net in $C_{usc}$ has an infimum in $C_{usc}$, which is given pointwise.
  2. (2)
    Each nondecreasing net in $C_{lsc}$ has a supremum in $C_{lsc}$, which is given pointwise.
Proof.
(1) Note that the empty set (or net) has infimum 00 in $C_{usc}$. Let $(g_{\alpha})_{\alpha}$ be a nonempty nonincreasing net in $C_{usc}$, and denote by $g:K\to[0,\infty)$ its pointwise infimum, which is given by $g(\rho)=\inf_{\alpha}g_{\alpha}(\rho)$ for each $\rho\in K$. Then $g$ is affine and upper semicontinuous by Lemma 2.8(1), whence $g\in C_{usc}$. It is then clear that $g$ is the infimum of the net $(g_{\alpha})_{\alpha}$ in $A(K)^{\infty}_{+}$, hence in $C_{usc}$.
(2) This is shown in a similar way using Lemma 2.8(2). ∎
Theorem 3.5.
Let $(V,V_{+},v)$ be a complete order unit space, with open cone $C:=V_{+}^{\circ}$.
  1. (1)
    Each $g\in C_{usc}$ is the pointwise infimum of the nonincreasing net $\{a\in C:g\prec a\}$.
  2. (2)
    Each $h\in C_{lsc}$ is the pointwise supremum of the nondecreasing net $\{b\in C:b\prec h\}$.
Proof.
(1) In Lemma 3.3 we have seen that $g$ is the pointwise infimum of the nonincreasing net $\mathfrak{A}:=\{a\in A(K):a\succ g\}$. Now for $a\in A(K)$ we have that $a\succ g$ and $g\geq 0$ imply $a\succ 0$. Therefore $\mathfrak{A}\subset C$ and the proof is complete.
(2) Let $\mathfrak{B}:=\{b\in A(K):b\prec h\}$. Applying Lemma 3.3(1) with $g=-h$ yields that for all $b_{1},b_{2}\in\mathfrak{B}$ there exists $b\in\mathbb{B}^{\prime}$ with $b\geq b_{i}$ for $i\in\{1,2\}$. Since $0\in\mathfrak{B}^{\prime}$ holds, the subnet $\mathfrak{B}^{\prime}:=\{b\in C:b\prec h\}=\{b\in A(K):0\prec b\prec h\}$ is cofinal in $\mathfrak{B}$. Therefore by part (2) of that theorem, $h(\rho)=\sup\{b(\rho):b\in\mathfrak{B}\}=\sup\{b(\rho):b\in\mathfrak{B}^{\prime}\}$ for every $\rho\in K$. ∎
Part (1) of the following lemma asserts that any nonincreasing net in $C_{usc}$ with infimum, say $g\in C_{usc}$, is cofinal in a suitable sense with respect to the net $\mathfrak{A}=\{x\in C:x\prec g\}$ considered in the previous. Part (2) contains the analogous cofinality assertion for a nondecreasing net in $C_{lsc}$. The lemma is a version of Dini’s theorem for semicontinuous functions.
Lemma 3.6.
Let $(V,V_{+},v)$ be a complete order unit space, with open cone $C:=V_{+}^{\circ}$.
  1. (1)
    Let $(g_{\alpha})_{\alpha}$ be any nonincreasing net in $C_{usc}$ with infimum $g\in C_{usc}$. Let $f\in C$ with $f\succ g$. Then there exist an index $\alpha_{0}$ and $\epsilon>0$ with $g_{\alpha_{0}}\prec f-\epsilon$.
  2. (2)
    Let $(h_{\alpha})_{\alpha}$ be any nondecreasing net in $C_{lsc}$ with supremum $h\in C_{lsc}$. Let $f\in C$ with $f\prec h$. Then there exist an index $\alpha_{0}$ and $\epsilon>0$ with $h_{\alpha_{0}}\succ f+\epsilon$.
Proof.
(1) Since $f$ is continuous, the functions $g_{\alpha}-f$ and $g-f$ are upper semicontinuous. From $g\prec f$ it follows by Lemma 2.9(1) that $\max(g-f)<0$, so there exists $\epsilon>0$ with $g-f\prec-\epsilon$. For every index $\alpha$ the set $U_{\alpha}:=\{\rho\in K:g_{\alpha}(\rho)-f(\rho)<-\epsilon\}$ is open since $U_{\alpha}=(g_{\alpha}-f)^{-1}((-\infty,\epsilon))$. We claim that $(U_{\alpha})_{\alpha}$ is an open cover of $K$. Indeed, because for each $\rho\in K$ we have that $\inf_{\alpha}(g_{\alpha}(\rho)-f(\rho))=g(\rho)-f(\rho)<-\epsilon$, there exists an index $\alpha$ such that $g_{\alpha}(\rho)-f(\rho)<-\epsilon$, that is, $\rho\in U_{\alpha}$. By compactness of $K$ we can find finitely many indices $\alpha_{1},\ldots,\alpha_{n}$ such that $K=\bigcup_{i=1}^{n}U_{\alpha_{i}}$. Since the net $(g_{\alpha})_{\alpha}$ is downward directed, there exists an index $\alpha_{0}$ such that for each $i\in\{1,2,\ldots,n\}$ we have $g_{\alpha_{i}}\geq g_{\alpha_{0}}$, in particular, $U_{\alpha_{i}}\subset U_{\alpha_{0}}$. It follows that $K=U_{\alpha_{0}}$. We conclude that $g_{\alpha_{0}}\prec f-\epsilon$, as desired.
(2) This is shown analogously to part (1). ∎
Lemma 3.7.
Let $K$ be the state space of the complete order unit space $(V,V_{+},v)$.
  1. (1)
    Let $g\colon K\to[-\infty,\infty)$ be affine and upper semicontinuous. Then there exists a pure state $\psi\in\delta_{e}K$ such that $\psi(g)=\sup g$.
  2. (2)
    Let $h\colon K\to(-\infty,\infty]$ be affine and lower semicontinuous. Then there exists a pure state $\psi\in\delta_{e}K$ such that $\psi(h)=\inf h$.
Proof.
(1) Let $F=\{\rho\in K:g(\rho)=\sup g\}$. Lemma 2.9 asserts that $g$ attains its supremum, so $F$ is nonempty. Moreover, $F$ is a face of $K$. Because $F=g^{-1}([\sup g,\infty))$ and $g$ is upper semicontinuous, $F$ is closed. We conclude that $F$ is a nonempty, closed face of $K$. The Krein–Milman Theorem 2.13 gives that $F$ is the closed convex hull of $\delta_{e}F$; in particular $\delta_{e}F\neq\emptyset$. Since $F$ is a face of $K$, we have $\delta_{e}K\cap F=\delta_{e}F\neq\emptyset,$ as was to be shown.
(2) Just like in part (1) one shows that $F=\{\rho\in K:h(\rho)=\inf h\}$ is a nonempty, closed face of $K$, which therefore contains an extremal point of $K$. ∎
Lemma 3.8.
Let $h,h^{\prime}\colon K\to(-\infty,\infty]$ be lower semicontinuous affine functions. If $h(\psi)\leq h^{\prime}(\psi)$ for each pure state $\psi\in\delta_{e}K$, then $h\leq h^{\prime}$.
Proof.
Theorem 3.5 furnishes an increasing net $(h_{\alpha})_{\alpha}$ of continuous affine functions on $K$ such that $h(\rho)=\sup_{\alpha}h_{\alpha}(\rho)$ for each $\rho\in K$. Now by replacing $h$ with an arbitrary member $h_{\alpha}$ of the net, which again satisfies $h_{\alpha}(\psi)\leq h^{\prime}(\psi)$ for each $\psi\in\delta_{e}K$, we may and do assume that $h$ is continuous.
The function $h_{0}=h^{\prime}-h$ is affine lower semicontinuous and satisfies $h_{0}(\psi)\geq 0$ for all $\psi\in\delta_{e}S$. It follows from Lemma 3.7(2) that $\inf h_{0}\geq 0$, whence $h^{\prime}\geq h$. ∎
We close this section by extending the gauge functions $L$ and $M$ defined in Definition 2.15 on $C$ to $A(K)^{\infty}_{+}$, and establishing some of their elementary properties.
Definition 3.9.
Let $x,y\in A(K)_{+}^{\infty}\setminus\{0,\infty\}$. We define the quantities
$m(y,x):=\sup\{\lambda\in[0,\infty):\lambda x\leq y\}\in[0,\infty),$ (29)
and
$M(x,y):=\inf\{\mu\in(0,\infty]:x\leq\mu y\}\in(0,\infty].$ (30)
Note that
$m(y,x)=M(x,y)^{-1}$ (31)
which is understood to include the assertion that $m(y,x)=0\iff M(x,y)=\infty$.
Lemma 3.10.
Let $x,y\in A(K)_{+}^{\infty}$ such that $x\succ 0$. If $\lambda\in[0,m(x,y))$ then $x-\lambda y\succ 0$.
Proof.
It is to be shown for every $\rho\in K$ that $x(\rho)-\lambda y(\rho)>0$. If $y(\rho)=0$ then we have $x(\rho)-\lambda y(\rho)=x(\rho)>0,$ since $x\succ 0$. Suppose next that $y(\rho)>0$. Then $m(x,y)y\leq x$ implies $x(\rho)\geq m(x,y)y(\rho)$. Therefore $x(\rho)-\lambda y(\rho)\geq(m(x,y)-\lambda)y(\rho)>0$, as $m(x,y)-\lambda>0$ by assumption. ∎
Lemma 3.11.
Let $x,y\in A(K)_{+}^{\infty}$ with $0\neq x\neq\infty$ and $y\prec\infty$. Let $\alpha\in(0,\infty)$, $\beta\in[0,\infty)$ and $\gamma\in[0,\alpha m(x,y))$. Then we have
$M(\alpha x+\beta y,x)$ $=\alpha+\beta M(y,x),$ (32)
$m(\alpha x+\beta y,x)$ $=\alpha+\beta m(y,x),$ (33)
$m(\alpha x-\gamma y,x)$ $=\alpha-\gamma M(y,x),$ (34)
$M(\alpha x-\gamma y,x)$ $=\alpha-\gamma m(y,x).$ (35)
Proof.
We content ourselves with proving formula (34), because the other formulae are obtained by similar arguments. Since $0\neq x\neq\infty$ we have $m(\alpha x,x)=\alpha$ so (34) holds if $\gamma=0$. Now if $0<\gamma<\alpha m(x,y)=m(\alpha x,y)$, then $\alpha x-\gamma y\in A(K)^{\infty}_{+}$ by Lemma 3.10. Now for $\lambda\in[0,\alpha/\gamma]$ we have
$\alpha-\gamma\lambda\leq m(\alpha x-\gamma y,x)\iff(\alpha-\gamma\lambda)x\leq\alpha x-\gamma y\iff y\leq\lambda x\iff\lambda\geq M(y,x).$
Using (the contrapositive of) this equivalence and the fact that $m(\alpha x-\gamma y,x)\leq m(\alpha x,x)=\alpha$ and $\alpha/\gamma>m(x,y)^{-1}=M(y,x),$ we conclude that
$m(\alpha x-\gamma y,x)$ $=\inf\{\alpha-\gamma\lambda:\lambda\in[0,\alpha/\gamma],\alpha-\gamma\lambda>m(\alpha x-\gamma y,x)\}$
$=\alpha-\gamma\sup\{\lambda\in[0,\alpha/\gamma]:\lambda<M(y,x)\}$
$=\alpha-\gamma M(y,x),$
which is this formula (34). ∎
3.2 Extension of the gauge-reversing bijection
The aim of this section is to show that each gauge-preserving or gauge-reversing bijection has an extension to the extended cones considered in Section 3.1 with the same properties.
Let $V$ and $W$ denote complete order unit spaces with open cones $C=V_{+}^{\circ}$ and $D=W_{+}^{\circ}$. Recall that the extended cones $C_{usc}$ and $C_{lsc}$ have been defined by (24) resp. (25), and we similarly define the extended cones $D_{usc}$ and $D_{lsc}$. Definitions 2.20 and 2.21 can naturally be generalized to maps between the extended cones, provided we impose an additional condition regarding the image of 00 or $\infty$. For example, a map $\Phi_{usc}\colon C_{usc}\to D_{lsc}$ is called gauge-reversing if $\Phi_{usc}(0)=\infty$ and $L(\Phi(x),\Phi(y))=m(x,y)$ for all $x,y\in C_{usc}\setminus\{0\}$.
Lemma 3.12.
For a bijection $\Phi$ from $C_{usc}$ to $D_{lsc}$, or from $C_{lsc}$ to $D_{usc}$, the following statements are equivalent:
  1. (1)
    $\Phi$ is gauge-reversing;
  2. (2)
    $\Phi$ is homogeneous of degree $-1$ and an order-anti-isomorphism.
Proof.
The proof, being analogous to Proposition 2.22, is omitted. ∎
Lemma 3.13.
For a bijection $\Psi$ from $C_{usc}$ to $D_{usc}$, or from $C_{lsc}$ to $D_{lsc}$, the following statements are equivalent:
  1. (1)
    $\Psi$ is gauge-preserving;
  2. (2)
    $\Psi$ is homogeneous of degree $1$ and an order-isomorphism.
Proof.
The proof, being analogous to Proposition 2.23, is omitted. ∎
Our aim now is to show that each gauge-reversing bijection $\Phi\colon C\to D$ extends to gauge-reversing bijections $\Phi_{usc}\colon C_{usc}\to D_{lsc}$ and $\Phi_{lsc}\colon C_{lsc}\to D_{usc}$. Observe that by criterion (2) of Lemma 3.12, these extensions are to be order-anti-isomorphisms, so are required to interchange suprema and infima. Because the members of the extended cones are limits of monotone nets in the original cones by Theorem 3.5, one arrives at the formulae (36) and (37) for $\Phi_{usc}$ resp. $\Phi_{lsc}$ in the following theorem.
Theorem 3.14.
Let $V$ and $W$ be complete order unit spaces, and let $\Phi\colon C\to D$ be a gauge-reversing bijection between their open cones $C=V^{\circ}_{+}$ and $D=W^{\circ}_{+}$. Then $\Phi$ has a unique extension to a gauge-reversing bijection
$\Phi_{usc}\colon C_{usc}\to D_{lsc}$
and to a gauge-reversing bijection
$\Phi_{lsc}\colon C_{lsc}\to D_{usc}.$
Proof.
We recall Proposition 2.22 that a map between (extended) cones is a gauge-reversing bijection if and only if it is homogeneous of degree $-1$ and an order-anti-isomorphism, i.e. an order-reversing bijection with an order-reversing inverse.
We define the extension $\Phi_{usc}:C_{usc}\to D_{lsc}$ as follows. Let $g\in C_{usc}$. By Theorem 3.5(1) the set $\{x\in C:x\succ g\}$ defines a nonincreasing net in $C$ with infimum $g$. It follows that $\{\Phi(x):x\succ g\}$ is a nondecreasing net in $D$, which has a supremum in $D_{lsc}$ by Lemma 3.4. We define
$\Phi_{usc}(g):=\sup\{\Phi(x):x\succ g\}\in D_{lsc}.$ (36)
Let us check that if $g\in C$ then $\Phi_{usc}(g)=\Phi(g)$. Since $\Phi$ is order-reversing, it is clear that $\Phi_{usc}(g)\leq\Phi(g)$. On the other hand, for every $\lambda>1$ we have $\lambda g\succ g$ and $\Phi(\lambda g)=\lambda^{-1}\Phi(g)$, so (36) gives $\Phi_{usc}(g)\geq\sup\{\lambda^{-1}\Phi(g):\lambda>1\}=\Phi(g)$. We conclude that $\Phi_{usc}(g)=\Phi(g)$.
Let us check that $\Phi_{usc}$ is order-reversing. Suppose that $g,g^{\prime}\in C_{usc}$ satisfy $g\leq g^{\prime}$. Then $\{x\in C:x\succ g^{\prime}\}\subseteq\{x\in C:x\succ g\}$, hence
$\Phi_{usc}(g^{\prime})=\sup\{\Phi(x):x\in C,x\succ g^{\prime}\}\leq\sup\{\Phi(x):x\in C,x\succ g\}=\Phi_{usc}(g).$
Let us check that $\Phi$ is homogeneous of degree $-1$. For every $\lambda\in(0,\infty)$ and $g\in C_{usc}$ we have $\{x\in C:x\succ\lambda g\}=\{\lambda x:x\in C,x\succ g\}$, so that
$\Phi_{usc}(\lambda g)$ $=\sup\{\Phi(\lambda x):x\in C,x\succ g\}=\sup\{\lambda^{-1}\Phi(x):x\in C,x\succ g\}$
$=\lambda^{-1}\sup\{\Phi(x):x\in C,x\succ g\}=\lambda^{-1}\Phi_{usc}(g).$
It remains to check that $\Phi_{usc}(0)=\infty$. For every $\epsilon>0$ we have $0\prec\epsilon v$, so that $\Phi(0)\geq\Phi(\epsilon v)=\epsilon^{-1}\Phi(v)$, from where $\Phi_{usc}(0)=\infty$.
Let us now show that $\Phi_{usc}$ is order-continuous with respect to nonincreasing nets. Let $(g_{\alpha})_{\alpha}$ be a nonincreasing net in $C_{usc}$, and let $g\in C_{usc}$ be its infimum. Since $\Phi_{usc}$ is order-reversing, the net $(\Phi_{usc}(g_{\alpha}))_{\alpha}$ in $D_{lsc}$ is nondecreasing. By definition we have $\Phi_{usc}(g_{\alpha})=\sup\{\Phi(x):x\in C,x\succ g_{\alpha}\}$, and $\Phi_{usc}(g)=\sup\{\Phi(x):x\in C,x\succ g\}$. As a consequence of Lemma 3.6 we obtain that $\{x\in C:x\succ g\}=\bigcup_{\alpha}\{x\in C:x\succ g_{\alpha}\}$. Now take the supremum to see that $\Phi_{usc}(g)=\sup_{\alpha}\Phi(g_{\alpha})$, as desired.
Similarly, using Theorem 3.5(2) that $\{x\in C:x\prec g\}$ is a nondecreasing net with supremum $g\in C_{lsc}$, we define
$\Phi_{lsc}(g)=\inf\{\Phi(x):x\in C,x\prec g\}.$ (37)
The verification that $\Phi_{lsc}$ is order-reversing, homogeneous of degree $-1$, and order-continuous with respect to nondecreasing nets goes along the same line.
Let us now show that $\Phi_{usc}\colon C_{usc}\to D_{lsc}$ is a bijection with inverse $(\Phi^{-1})_{lsc}\colon D_{lsc}\to C_{usc}$. Let $g\in C_{usc}$, and choose some nonincreasing net $(g_{\alpha})$ in $C$ with infimum $g$. Because $\Phi_{usc}$ is order-continuous with respect to nonincreasing nets, $\Phi(g_{\alpha})$ is a nondecreasing net in $D$ with supremum $\Phi_{usc}(g)$ in $D_{usc}$. Similarly, using that $(\Phi^{-1})_{lsc}$ is order-continuous with respect to nondecreasing nets, we obtain that $(\Phi^{-1})_{lsc}(\Phi_{usc}(g))=\inf_{\alpha}\Phi^{-1}(\Phi(g_{\alpha}))=\inf_{\alpha}g_{\alpha}=g$. This shows that $(\Phi^{-1})_{lsc}\circ\Phi_{usc}=\operatorname{Id}_{C_{usc}}$; the proof that $\Phi_{usc}\circ(\Phi^{-1})_{lsc}=\operatorname{Id}_{D_{lsc}}$ is similar. Since $\Phi_{usc}$ and $(\Phi^{-1})_{lsc}$ are mutually inverse order-reversing bijections, they are order-anti-isomorphisms, and also homogeneous of degree $-1$. We conclude by Proposition 2.22 that $\Phi_{usc}$ and $(\Phi^{-1})_{lsc}$ are gauge-reversing bijections. Considering the gauge-reversing map $\Phi^{-1}$ instead, we obtain that $\Phi_{lsc}$ and $(\Phi^{-1})_{usc}$ are mutually inverse bijections.
Since an order-anti-isomorphism sends the infimum of a set to the supremum of the image of that set, any gauge-reversing bijection $\Phi_{usc}\colon C_{usc}\to D_{lsc}$ extending $\Phi$ is necessarily given by (36). This shows the asserted uniqueness of $\Phi_{usc}$. Similarly, the gauge-reversing bijective extension $\Phi_{lsc}$ of $\Phi$ is uniquely determined by (37). ∎
Now suppose $\Psi\colon C\to D$ that is a gauge-preserving bijection. By the same method as in Theorem 3.14, one constructs extensions $\Psi_{usc}\colon C_{usc}\to D_{usc}$ and $\Psi_{lsc}\colon C_{lsc}\to D_{lsc}$.
Proposition 3.15.
Let $V$ and $W$ be complete order unit spaces, and let $\Psi\colon C\to D$ be a gauge-preversing bijection between their open cones $C=V^{\circ}_{+}$ and $D=W^{\circ}_{+}$. Then $\Psi$ has a unique extension to a gauge-preversing bijection
$\Psi_{usc}\colon C_{usc}\to D_{usc}\quad\text{ and }$
and to a gauge-preserving bijection
$\Psi_{lsc}\colon C_{lsc}\to D_{lsc}.$
Proof.
Modify the proof of Theorem 3.14 appropriately. ∎
3.3 Strong atomicity
Definition 3.16.
Let $(X,X_{+})$ be an ordered vector space, and let $E\subset X_{+}$ be a subcone containing 00. An extremal vector of $E$ is a nonzero vector $p\in E$ such that $\{x\in E:x\leq p\}=\{tp:0\leq t\leq 1\}$. The set of extremal vectors of $E$ is denoted $\operatorname{ext}E$. We say that the cone $E$ is strongly atomic if for each $x\in E$ one has
$x=\sup\{p\in\operatorname{ext}E\colon p\leq x\}\text{ in }E.$
We continue with our assumption that $(V,V_{+},v)$ and $(W,W_{+},w)$ be complete order unit spaces, with state spaces $K$ resp. $S$, and open cones $C=V^{\circ}_{+}$ and $D=W^{\circ}_{+}$. The set of extreme vectors $p$ in $C_{usc}$ normalized so that $M(p,v)=1$, or equivalently so that $m(v,p)=1$, is denoted by
$\operatorname{ext}_{v}C_{usc}:=\{p\in\operatorname{ext}C_{usc}:M(p,v)=1\}.$
Lemma 3.17.
Let $(V,V_{+},v)$ be an order unit space. Then the following are equivalent:
  1. (1)
    the cone $C_{usc}$ is strongly atomic;
  2. (2)
    for each $g\in C_{usc}$ we have
    $g=\sup\{m(g,p)p:p\in\operatorname{ext}_{v}C_{usc}\}$ (38)
  3. (3)
    for all $g,g^{\prime}\in C_{usc}$ we have
    $m(g,p)\leq m(g^{\prime},p)\text{ for each }p\in\operatorname{ext}_{v}C_{usc}\iff g\leq g^{\prime}.$ (39)
Proof.
As to (1) $\iff$ (2), note that $\{r\in\operatorname{ext}C_{usc}:r\leq g\}=\{m(g,p)p:p\in\operatorname{ext}_{v}C_{usc}\}.$ The equivalence (2) $\iff$ (3) holds because for each $p\in\operatorname{ext}_{v}C_{usc}$ we have that $m(g^{\prime},p)\geq m(g,p)$ if and only if $m(g,p)p\leq g^{\prime}$. ∎
Recall that we defined $\delta_{e}S$ as the set of pure states in $S$, and for each $\psi\in\delta_{e}S$ the function $I^{\infty}_{\{\psi\}}$ by $I^{\infty}_{\{\psi\}}(\psi)=1$ and $I^{\infty}_{\{\psi\}}(\rho)=\infty$ if $\rho\neq\psi$, see Example 3.1.
Proposition 3.18.
Let $\Phi\colon C\to D$ be a gauge-reversing bijection such that $\Phi(v)=w$. Then we have a bijection
$\operatorname{ext}_{v}C_{usc}$ $\to\delta_{e}S,$ (40)
$p$ $\mapsto\psi,$ (41)
where $\psi$ is determined by $I^{\infty}_{\{\psi\}}=\Phi_{usc}(p)$. Moreover, for all $g\in C_{usc}$ we have
$M(p,g)=\Phi_{usc}(g)(\psi).$ (42)
Proof.
Let $p\in C_{usc}$ and $f\in D_{lsc}$ be such that $f=\Phi_{usc}(p)$, hence $p=(\Phi^{-1})_{lsc}(f)$. We need only show that $p\in\operatorname{ext}_{v}C_{usc}$ if and only if $f=I^{\infty}_{\{\psi\}}$ for some $\psi\in\delta_{e}S$. Since the bijection $\Phi_{usc}$ is an order-anti-isomorphism and homogeneous of degree $-1$, it restricts to bijections
[…]
Moreover, we have $m(v,p)=m(f,w)=\inf f$, so we may assume that $m(v,p)=1=\inf f$. Then $p\in\operatorname{ext}C_{usc}$ if and only if the left inclusion is an equality, which happens only if the right inclusion is an equality. We will finish the proof by showing that we have equality on the right if and only if $f=I^{\infty}_{\{\psi\}}$ for some $\psi\in\delta_{e}S$.
The if direction being clear, suppose we have equality on the right. By Lemma 3.7 there exists $\psi\in\delta_{e}S$ with $f(\psi)=\inf f=1$. Then we have $I^{\infty}_{\psi}\geq f$, hence there exists $\lambda\in[1,\infty]$ with $I^{\infty}_{\psi}=\lambda f$. Then $\lambda=\inf(\lambda f)=\inf I^{\infty}_{\psi}=1$, so $f=I^{\infty}_{\psi}$ as desired.
Finally, we turn to proving (42). Note that for each $h\in D_{lsc}$ and $\psi\in\delta_{e}S$ we have that $h(\psi)=\inf\{\lambda>0:h\leq\lambda I^{\infty}_{\{\psi\}}\}=M(h,I^{\infty}_{\{\psi\}})$. Therefore if $p\in\operatorname{ext}_{v}C_{usc}$ and $g\in C_{usc}$ are such that $\Phi_{usc}(p)=I^{\infty}_{\{\psi\}}$ and $\Psi_{usc}(g)=h$, then we have
$\Phi_{usc}(g)(\psi)=h(\psi)=M(h,I^{\infty}_{\{\psi\}})=M(\Phi_{usc}(g),\Phi_{usc}(p))=M(p,g),$
where in the last step we used that $\Phi_{usc}$ is gauge-reversing. ∎
Theorem 3.19.
Let $\Phi\colon C\to D$ be a gauge-reversing bijection. Then $\Phi$ is convex, i.e. for all $x_{1},x_{2}\in C$ and $\lambda\in[0,1]$ one has
$\Phi((1-\lambda)x_{1}+\lambda x_{2})\leq(1-\lambda)\Phi(x_{1})+\lambda\Phi(x_{2}).$ (43)
Proof.
Let $\psi\in\delta_{e}S$ be a pure state of $W$. We will show that the function $\psi\circ\Phi\colon C\to(0,\infty)$ is convex. Proposition 3.18 grants us an extremal vector $p\in\operatorname{ext}_{v}(C_{usc})$ such that $\psi(\Phi(z))=M(p,z)=m(z,p)^{-1}$ for every $z\in C$. It suffices to show that the map $m(\cdot,p)\colon C\to(0,\infty)$ is concave, because the inverse of a strictly positive concave function is convex.
We claim that $m(\cdot,p)$ is homogeneous of degree $1$ and superadditive. Plainly, for all $\lambda>0$ and $z\in C$ we have $m(\lambda z,p)=\lambda m(z,p)$. Let now $z_{1},z_{2}\in C$. Then $m(z_{1},p)p\leq z_{1}$ and $m(z_{2},p)p\leq z_{2}$ hold, so $(m(z_{1},p)+m(z_{2},p))p\leq z_{1}+z_{2}$, whence $m(z_{1},p)+m(z_{2},p)\leq m(z_{1}+z_{2},p)$. This proves our claim. It follows that the function $m(\cdot,p)$ is concave, since
$m((1-\lambda)x_{1}+\lambda x_{2},p)\geq m(1-\lambda)x_{1},p)+m(\lambda x_{2},p)=(1-\lambda)m(x_{1},p)+\lambda m(x_{2},p).$
Consequently, $\psi\circ\Phi=m(\cdot,p)^{-1}$ is convex. Since $\psi\in\delta_{e}S$ was arbitrary, we conclude that $\Phi$ is convex by Lemma 3.8. ∎
Theorem 3.20.
Let $(V,V_{+},v)$ and $(W,W_{+},w)$ be complete order unit spaces such that there exists a gauge-reversing bijection $\Phi\colon C\to D$. Then the cone $C_{usc}$ is strongly atomic.
Proof.
Let us show that condition (3) of Lemma 3.17 is satisfied. Let $g,g^{\prime}\in C_{usc}$ be such that $m(g,p)\leq m(g^{\prime},p)$, equivalently $M(p,g)\geq M(p,g^{\prime})$, for each $p\in\operatorname{ext}_{v}C_{usc}$. Set $h=\Phi_{usc}(g)$ and $h^{\prime}=\Phi_{usc}(g^{\prime})$. Let $\psi\in\delta_{e}S$ be arbitrary. By Proposition 3.18 there exists a unique $p\in\operatorname{ext}_{v}C_{usc}$ such that $\Phi_{usc}(p)=I^{\infty}_{\{\psi\}}$. Moreover, according to (42) we have that $M(p,g)=h(\psi)$ and $M(p,g^{\prime})=h^{\prime}(\psi)$. It follows that $h(\psi)=M(p,g)\geq M(p,g^{\prime})=h^{\prime}(\psi).$ Now Lemma 3.8 implies that $h\geq h^{\prime}$. Since $\Phi_{usc}$ is an order-anti-isomorphism, we conclude that $g\leq g^{\prime}$. ∎
Corollary 3.21.
Let $y,z\in C_{usc}$. Then
$\lVert y-z\rVert_{v}\leq\sup\{|m(y,p)-m(z,p)|:p\in\operatorname{ext}_{v}C_{usc}\}.$
Proof.
Denote the supremum on the right by $\nu$. We claim that $y\leq z+\nu v$. Indeed, for each $p\in\operatorname{ext}_{v}C_{usc}$ we have $m(z+\nu v,p)\geq m(z,p)+m(\nu v,p)=m(z,p)+\nu\geq m(y,p)$, so by Theorem 3.20 it follows that $z+\nu v\geq y$. Now $y+\nu v\geq z$ follows by symmetry, whence $-\nu v\leq y-z\leq\nu v$. This shows that $\lVert y-z\rVert_{v}\leq\nu$, as desired. ∎
4 Consequences of strong atomicity
In Theorem 3.20 of the previous section we have established that the extended cone $C_{usc}$ is strongly atomic. Each of the three subsections of this section derives a distinct consequence of strong atomicity of $C_{usc}$. In the first subsection we prove that the gauge-reversing bijection $\Phi\colon C\to D$ is Gateaux-differentiable in positive directions, and that the negative of the Gateaux derivative at a point $x\in C$ extends to an isomorphism $(V,V_{+},x)\cong(W,W_{+},\Phi(x))$. In the second subsection, we give an alternative proof of Theorem 2.24, stating that a gauge-preserving bijection $\Psi\colon C\to D$ extends to a linear isomorphism between $V$ and $W$, under the additional assumption of strong atomicity of $C_{usc}$. Finally, in the third subsection, we prove a strong form of Hua’s identity, which provides an explicit formula for $D\Phi$ in terms of $\Phi$.
4.1 Gateaux-differentiability
In this section we will show that the gauge-reversing bijection $\Phi_{usc}\colon C_{usc}\to D_{lsc}$ extending $\Phi$ has a (one-sided) Gateaux derivative at every point $x\in C$ in each direction $y\in C_{usc}$ (see Section 2.1 for a discussion of Gateaux-differentiability). This Gateaux-derivative, denoted $D\Phi(x)(y)$, belongs to $-D_{usc}$ since $\Phi_{usc}$ is order-reversing. We show that the map $-D\Phi(x):C_{usc}\to D_{usc}$ extends to an isomorphism $(V,V_{+},x)\cong(W,W_{+},\Phi(x))$, culminating in Theorem 4.7.
We will first prove Gateaux-differentiability in the direction of an extremal vector of $C_{usc}$, by adopting the arguments [LRW25, Lemmata 2.1 and 6.1] to the context of extended cones.
Lemma 4.1.
Let $x\in C$ and $y\in C_{usc}$ with $y>x$. Then the following are equivalent:
  1. (1)
    we have that $y-x\in C_{usc}$ is an extremal vector of $C_{usc}$;
  2. (2)
    we have $\{z\in C_{usc}:x\leq z\leq y\}=\{(1-t)x+ty:0\leq t\leq 1\}$;
  3. (3)
    the set $\{z\in C_{usc}:x\leq z\leq y\}$ is totally ordered.
Proof.
The validity of each of the $3$ statements is unaltered if we subtract $x$ from $x$ and $y$, so we may and do assume that $x=0$. Then (1) $\iff$ (2) holds by Definition 3.16, while (2) $\implies$ (3) is plain from $y\geq 0$. As to (3) $\implies$ (2), note that the containment $\supseteq$ in (2) always holds.
Now assume $\{z\in C_{usc}:0\leq z\leq y\}$ is totally ordered, and let $z$ be one of its members. We finish the proof by showing that $z=\lambda y$ for some $\lambda\in[0,1]$. For every $\mu\in[0,1]$ either $\mu y\leq z$ or $\mu y\geq z$. Let $\lambda=\sup\{\mu\in[0,1]:\mu y\leq z\}=\inf\{\mu\in[0,1]:\mu y\geq z\}$. For each $\epsilon>0$ we have $(\lambda-\epsilon)y\leq z\leq(\lambda+\epsilon)y$, or $-\epsilon y\leq z-\lambda y\leq\epsilon y$, and since $(V,V_{+})$ is Archimedean this implies $z-\lambda y=0$. We conclude that $z\in\{\lambda y:\lambda\in[0,1]\}$ as desired. ∎
Lemma 4.2.
Let $x\in C$ and $y\in C_{lsc}$ with $y<x$. Then the following are equivalent:
  1. (1)
    we have that $x-y\in C_{usc}$ is an extremal vector of $C_{usc}$;
  2. (2)
    we have $\{z\in C_{lsc}:y\leq z\leq x\}=\{(1-t)x+ty:0\leq t\leq 1\}$;
  3. (3)
    the set $\{z\in C_{lsc}:y\leq z\leq x\}$ is totally ordered.
Proof.
This is deduced from Lemma 4.1 using the order-anti-isomorphism $z\mapsto x-z$ between $\{z\in C_{lsc}:y\leq z\leq x\}$ and $\{z\in C_{usc}:0\leq z\leq x-y\}$. ∎
Since condition (3) in Lemma 4.1 and Lemma 4.2 is purely order-theoretic, it follows that a gauge-reversing bijection sends extremal rays to extremal rays. A calculation of gauge-functions then yields the following more precise results.
Theorem 4.3.
Let $\Phi\colon C\to D$ be a gauge-reversing bijection between the open cones of complete order unit spaces.
  1. (1)
    Let $x\in C$ and $p\in\operatorname{ext}(C_{usc})$ with $m(x,p)=1$. Then there exists $q\in\operatorname{ext}(D_{usc})$ with $m(\Phi(x),q)=1$ such that for all $\lambda\in(0,\infty)$ one has
    $\Phi_{usc}(x+\lambda p)=\Phi(x)-\frac{\lambda}{\lambda+1}q\in D_{lsc}.$ (44)
  2. (2)
    Let $z\in D$ and $q\in\operatorname{ext}(D_{usc})$ with $m(z,q)=1$. Then there exists $p\in\operatorname{ext}(C_{usc})$ with $m(\Phi^{-1}(z),p)=1$ such that for all $\lambda\in(-1,0)$ one has
    $(\Phi^{-1})_{lsc}(z+\lambda q)=\Phi^{-1}(z)-\frac{\lambda}{\lambda+1}p\in C_{usc}.$ (45)
Proof.
Fix some large $\kappa>0$ for the moment. Since $\kappa p$ is an extremal vector, the order interval $\{y\in C_{usc}:x\leq y\leq x+\kappa p\}$ is totally ordered by Lemma 4.1. Since $\Phi_{usc}$ is an order-anti-isomorphism, it follows that
$\Phi_{usc}(\{y\in C_{usc}:x\leq y\leq x+\kappa p\})=\{z\in D_{lsc}:\Phi_{usc}(x+\kappa p)\leq z\leq\Phi(x)\}$
is totally ordered as well. This means that condition (3) of Lemma 4.2 is satisfied, and so the other two conditions also hold. First, conditions (1) gives that $\Phi(x)-\Phi_{usc}(x+\kappa p)$ is an extremal vector of $D_{usc}$, hence a multiple of an extremal vector $q$ with $M(q,\Phi(x))=1$. Secondly, condition (2) implies that for each $0\leq\lambda\leq\kappa$ there exists $t(\lambda)\leq 0$ with $\Phi(x+\lambda p)=\Phi(x)+t(\lambda)q$. To calculate $t(\lambda)$, use equations (32) and (34) as well as the fact that $m(x,p)=1=m(\Phi(x),q)$ to see that
$\frac{1}{1+\lambda}=m(x,x+\lambda p)=m(\Phi(x)+t(\lambda)q,\Phi(x))=1+t(\lambda).$
It follows that $t(\lambda)=1/(1+t)-1=-t/(t+1)$. Since increasing $\kappa$ does not change $q$, we conclude that (44) holds for all $\lambda\geq 0$. Part (2) is shown in a similar fashion. ∎
Proposition 4.4.
Let $x\in C$. There exists a homogeneous of degree $1$, order-reversing map
$D\Phi_{usc}(x)\colon C_{usc}\to-D_{usc},$
such that $D\Phi_{usc}(x)(y)$ is the Gateaux-derivative of $\Phi_{usc}$ at $x$ in the direction $y$ for each $y\in C_{usc}$. Moreover, $D\Phi_{usc}(x)$ maps $V_{+}$ into $-W_{+}$, sends $x$ to $-\Phi(x)$, and maps $C$ into $-D$.
Proof.
There is no loss of generality in assuming $x=v$ and $\Phi(x)=w$, cf. Remark 3.2. Fix $y\in C_{usc}$, and define $f\colon(0,\infty)\to D_{usc}$ by the difference quotient
$f(\mu):=\mu^{-1}(w-\Phi_{usc}(v+\mu y)).$
Note here that $v+\mu y\geq v$ implies $w\geq\Phi_{usc}(v+\mu y)$, so that $\mu^{-1}(w-\Phi_{usc}(v+\mu y))\in D_{usc}$. We will show there exists $f(0)\in D_{usc}$ such that
$\lim_{\mu\downarrow 0}\lVert f(\mu)-f(0)\rVert_{w}=0.$ (46)
Let $q\in\operatorname{ext}_{w}D_{usc}$ and use Theorem 4.3(2) to obtain $p\in\operatorname{ext}_{v}C_{usc}$ such that (45) holds with $z=w$ and $\Phi^{-1}(z)=v$. Let $\mu,\nu\in(0,\infty)$. We claim that
$m(f(\mu),q)=\frac{m(y,p)}{1+\mu m(y,p)}.$ (47)
For $\lambda\geq 0$ the following are equivalent:
$\lambda\leq\mu m(y,p)\iff\lambda p\leq\mu y$ $\iff v+\lambda p\leq v+\mu y$
$\iff\Phi(v+\mu y)\leq\Phi(v+\lambda p)=w-\frac{\lambda}{1+\lambda}q$
$\iff\frac{\lambda}{1+\lambda}q\leq w-\Phi(v+\mu y).$
It follows that
$m(\mu^{-1}(w-\Phi(v+\mu y)),q)=\mu^{-1}\cdot\frac{\mu m(y,p)}{1+\mu m(y,p)}=\frac{m(y,p)}{1+\mu m(y,p)},$
whence our claim follows. A little calculation shows that
$m(f(\mu),q)-m(f(\nu),q)=\frac{m(y,p)}{1+\mu m(y,p)}-\frac{m(y,p)}{1+\nu m(y,p)}=\frac{(\nu-\mu)m(y,p)^{2}}{(1+\mu m(y,p))(1+\nu m(y,p))}.$ (48)
Since $0\leq m(y,p)\leq\lVert y\rVert_{v}$, we conclude that
$\sup\{|m(f(\mu),q)-m(f(\nu),q)|:q\in\operatorname{ext}_{w}D_{usc}\}\leq\lVert y\rVert_{v}^{2}|\mu-\nu|.$
By Corollary 3.21 it follows that
$\lVert f(\mu)-f(\nu)\rVert_{w}\leq\lVert y\rVert_{v}^{2}|\mu-\nu|.$ (49)
Since each $\lVert\cdot\rVert_{w}$-Cauchy net in $D_{usc}$ converges in $D_{usc}$ by Lemma 2.8(1), it follows that there exists $f(0)\in D_{usc}$ satisfying (46). If $y\in V_{+}$, then $f(\mu)\in W_{+}$ for all $\mu>0$ so $f(0)\in W_{+}$ because $(W_{+},\lVert\cdot\rVert_{w})$ is complete. Setting $D\Phi_{usc}(v)(y)=-f(0)\in-D_{usc}$ and letting $\nu\downarrow 0$ in (49) gives
$\lVert\mu^{-1}(\Phi_{usc}(v+\mu y)-\Phi(v))-D\Phi_{usc}(v)(y)\rVert_{w}=\lVert f(\mu)-f(0)\rVert_{w}\leq\mu\lVert y\rVert_{v}^{2}.$ (50)
We conclude that $D\Phi_{usc}(v)(y)$ is the Gateaux-derivative of $\Phi_{usc}$ at $v$ in the direction $y$.
Now let $x\in C$ be arbitrary. By a general principle of Gateaux-derivatives, $D\Phi_{usc}(x)$ is homogeneous of degree $1$. It reverses the order since $\Phi_{usc}$ does. Since $\Phi$ is homogeneous of degree $-1$, one easily sees that $D\Phi_{usc}(x)(x)=-\Phi(x)$. This implies by homogeneity that $D\Phi_{usc}(x)(\delta x)=-\delta\Phi(x)$ for each $\delta>0$. Because we have shown that $D\Phi_{usc}(x)$ maps $V_{+}$ into $-W_{+}$ and is order-reversing, it follows that $C=\bigcup_{\delta>0}\{y\in V_{+}:y\geq\delta x\}$ is mapped into $-D=\bigcup_{\delta>0}\{z\in-W_{+}:z\leq-\delta\Phi(x)\}$ by $D\Phi_{usc}(x)$. ∎
Proposition 4.5.
Let $x\in C$. There exists a homogeneous order-reversing map
$D\Phi_{lsc}(x)\colon-C_{usc}\to D_{usc},$
such that $D\Phi_{lsc}(x)(-y)$ is the Gateaux-derivative of $\Phi_{lsc}$ at $x$ in the direction $-y$ for each $y\in C_{usc}$. Moreover, $D\Phi_{lsc}(x)$ maps $-V_{+}$ inside $W_{+}$, sends $-x$ to $\Phi(x)$, and maps $-C$ into $D$.
Proof.
The proof being very similar to that of Proposition 4.4, we only indicate which adaptions ought to be made. Again we may assume that $x=v$ and $\Phi(x)=w$. Let $y\in C_{usc}$ and define $f\colon(-\lVert y\rVert_{v},0)\to D_{usc}$ by
$f(\mu)=\mu^{-1}(w-\Phi_{lsc}(v+\mu y)).$
Now one establishes for $-\max\{1,\lVert y\rVert_{v}\}^{-1}<\mu<\nu<0$ that (47) and (48) hold, and deduces from Corollary 3.21 that
$\lVert f(\nu)-f(\mu)\rVert_{w}\leq\frac{(\nu-\mu)\lVert y\rVert_{v}^{2}}{(1+\mu\lVert y\rVert_{v})^{2}}.$
The proof is concluded in the same way as the previous proposition. ∎
Lemma 4.6.
The maps $D\Phi_{usc}(x)\colon C_{usc}\to-D_{usc}$ and $D(\Phi^{-1})_{lsc}(\Phi(x))\colon-D_{usc}\to C_{usc}$ are mutually inverse for every $x\in C$.
Proof.
The maps $\Phi_{usc}\colon C_{usc}\to D_{lsc}$ and $(\Phi^{-1})_{lsc}\colon D_{lsc}\to C_{usc}$ are mutually inverse by Theorem 3.14. As we have seen in Theorem 2.29, both maps are locally Lipschitz continuous, so the chain rule for Gateaux-derivatives, Lemma 2.2, holds:
$D(\Phi^{-1})_{lsc}(\Phi(x))\circ D\Phi_{usc}(x)=\operatorname{id}_{C_{usc}}\text{ and }D\Phi_{usc}(x)\circ D(\Phi^{-1})_{lsc}(\Phi(x))=\operatorname{id}_{-D_{usc}}.$
Theorem 4.7.
For each $x\in C$ there exists a unique isomorphism of order unit spaces
$-D\Phi(x)\colon(V,V_{+},x)\to(W,W_{+},\Phi(x))$ (51)
such that for all $y\in V_{+}$ one has
$\lVert\Phi(x+y)-\Phi(x)-D\Phi(x)(y)\rVert_{\Phi(x)}\leq\lVert y\rVert^{2}_{x}.$ (52)
Proof.
Propositions 4.4 and 4.5 and Lemma 4.6 yield that the map $-D\Phi_{usc}(x)\colon C_{usc}\to D_{usc}$ is an order-isomorphism and homogeneous of degree $1$, hence a gauge-preserving bijection by Proposition 2.22. By the same propositions, $-D\Phi_{usc}(x)$ restricts to a gauge-preserving bijection from $C$ into $D$, which we denote by $-D\Phi(x)\colon C\to D$. By Theorem 2.24 the map $-D\Phi(x)\colon C\to D$ extends to a linear isomorphism also denoted by $-D\Phi(x)\colon V\to W.$ Since $D\Phi(x)(x)=-\Phi(x)$ holds by Proposition 4.4, we have that $-D\Phi(x)\colon(V,V_{+},x)\to(W,W_{+},\Phi(x))$ is an isomorphism of order unit spaces.
Taking $\mu=1$ in (50) we obtain that for all $y\in V_{+}$ we have
$\lVert\Phi(v+y)-\Phi(v)-D\Phi(v)(y)\rVert_{w}\leq\lVert y\rVert_{v}^{2}.$
Since the order unit of $(V,V_{+})$ may be chosen at will, the last formula holds with $v$ replaced by an arbitrary element $x$ of $C$ and $w$ replaced by $\Phi(x)$. ∎
Remark 4.8.
Theorem 4.7 yields a proof of Theorem 1.3 in the case that both order unit spaces are complete. We will dispense with this completeness assumption in Section 5.8.
4.2 Noll-Schäfer theorem for strongly atomic cones
Let $(V,V_{+},v)$ and $(W,W_{+},w)$ be complete order unit spaces. By a theorem of Noll and Schäfer, [NS77, Theorem B], if $\Psi\colon C\to D$ is a gauge-preserving bijection with $\Psi(v)=w$, then $\Psi$ extends to an isomorphism of order unit spaces $(V,V_{+},v)\cong(W,W_{+},w)$. To make our exposition more self-contained, we will give an independent proof of this theorem under the assumption that the cone $C_{usc}$ be strongly atomic. Note that by Theorem 3.20 the latter condition is satisfied if there exists a gauge-reversing map $\Phi\colon C\to D$.
Theorem 4.9.
Let $(V,V_{+},v)$ and $(W,W_{+},w)$ be complete order unit spaces, and suppose that $C_{usc}$ is strongly atomic. Let $\Psi\colon C\to D$ be a gauge-preserving bijection with $\Psi(v)=w$. Then $\Psi$ extends to an isomorphism of order unit spaces $(V,V_{+},v)\cong(W,W_{+},w)$.
Proof.
Let $\Psi_{usc}\colon C_{usc}\to D_{usc}$ be the gauge-preserving bijection extending $\Psi$ that has been constructed in Proposition 3.15. We will show that $\Psi_{usc}(x+y)=\Psi_{usc}(x)+\Psi_{usc}(x)$ for all $x,y\in C_{usc}$.
First assume that $x\in C$. For each $p\in\operatorname{ext}C_{usc}$ there exists $q\in\operatorname{ext}D_{usc}$ such that for each $\lambda\geq 0$ one has $\Psi_{usc}(x+\lambda p)=\Psi(x)+\lambda q$. Then for $\lambda\geq 0$ and $\mu>0$, we have
$x+\lambda p\leq x+\mu y\iff\Psi(x)+\lambda q=\Psi_{usc}(x+\lambda p)\leq\Psi_{usc}(x+\mu y).$
It follows that
$m(y,p)=m(\mu^{-1}(\Psi_{usc}(x+\mu y)-\Psi(x)),q).$
Since the left-hand side does not depend on $\mu$, neither does the right-hand side. By Corollary 3.21 we conclude that for all $\mu>0$ one has
$\mu^{-1}(\Psi_{usc}(x+\mu y)-\Psi(x))=\Psi_{usc}(x+\mu y)-\Psi(x).$ (53)
Now if $x\in C_{usc}$ is arbitrary, we may take a nonincreasing net $(x_{\alpha})_{\alpha}$ in $C$ with infimum $x$ to see that (53) holds for all $x,y\in C_{usc}$ and $\mu>0$. We obtain that
$\Psi_{usc}(x+y)$ $=\mu^{-1}\Psi_{usc}(\mu x+\mu y)$
$=\mu^{-1}(\Psi_{usc}(\mu x+\mu y)-\Psi_{usc}(\mu x))+\mu^{-1}\Psi_{usc}(\mu x)$
$=\Psi_{usc}(\mu x+y)-\Psi_{usc}(\mu x)+\Psi_{usc}(x).$
Taking the infimum over all $\mu>0$ gives that
$\Psi_{usc}(x+y)=\Psi_{usc}(x)+\Psi_{usc}(y).$
We conclude that $\Psi_{usc}$ is additive on $C_{usc}$, so in particular $\Psi$ is additive on $V_{+}$.
Since $V=V_{+}-V_{+}$ there exists a unique linear map extending $\Psi$, which we denote $\Psi\colon V\to W$ as well. Similarly, $\Psi^{-1}\colon D\to C$ extends to $\Psi^{-1}\colon W\to V$. It is clear that $\Psi^{-1}\circ\Psi=\operatorname{id}_{V}$ and $\Psi\circ\Psi^{-1}=\operatorname{id}_{W}$, so $\Psi$ is a linear bijection. Moreover, $\Psi(C)=D$ implies that $\Psi(V_{+})=W_{+}$. We conclude that $\Psi\colon(V,V_{+},v)\to(W,W_{+},w)$ is an isomorphism of order unit spaces. ∎
4.3 Hua’s identity
The principal result of this subsection is that every gauge-reversing bijection $\Phi$ automatically satisfies Hua’s identity. This identity will play a central role in all subsequent subsections. As a corollary we obtain a formula for the Gateaux-derivative $D\Phi$ of $\Phi$ solely in terms of $\Phi$ itself, which will be instrumental in establishing higher-order smoothness properties of $\Phi$.
Theorem 4.10 (Hua’s identity).
Let $\Phi\colon C\to D$ be a gauge-reversing bijection between the open cones of complete order unit spaces $(V,V_{+},v)$ resp. $(W,W_{+},w)$. Then for all $x,y\in C$ we have
$\Phi(x)-\Phi(x+y)=-D\Phi(x)(\Phi^{-1}(\Phi(x)+\Phi(y))).$ (54)
Proof.
In view of Remark 3.2, we may and do assume that $x=v$ and $\Phi(x)=w$, so our objective becomes to show the equation
$w-\Phi(v+y)=-D\Phi(v)(\Phi^{-1}(w+\Phi(y))).$
Both members reside in $D$, since $w-\Phi(v+y)\geq w-\Phi(v+m(y,v)v)=(1-(1+m(y,v))^{-1})w$ and because $-D\Phi(v):(V,V_{+},v)\to(W,W_{+},w)$ is an isomorphism of order unit spaces by Theorem 4.7.
By Theorem 3.20 it suffices to prove for each $q\in\operatorname{ext}_{w}D_{usc}$ that
$m(w-\Phi(v+y),q)=m({-D\Phi(v)}(\Phi^{-1}(w+\Phi(y))),q).$ (55)
Let $p\in\operatorname{ext}_{v}C_{usc}$ and $q\in\operatorname{ext}_{w}D_{usc}$ be a pair of extremal vectors corresponding to one another as in Theorem 4.3(1), i.e. such that for each $\lambda>0$ we have
$\Phi(v+\lambda p)=w-\frac{\lambda}{\lambda+1}q.$ (56)
To calculate the left-hand side of (55), observe that for each $\lambda>0$ it holds that
$\lambda p\leq y\iff v+\lambda p\leq v+y\iff w-\frac{\lambda}{\lambda+1}q\geq\Phi(v+y)\iff\frac{\lambda}{\lambda+1}q\leq w-\Phi(v+y).$
Because we defined $m(y,p)=\sup\{\lambda>0:\lambda p\leq y\}$, this shows that
$m(w-\Phi(v+y),q)=\frac{m(y,p)}{m(y,p)+1}.$
We now turn to the right-hand side of (55). One sees from (56) that the Gateaux-derivative of $\Phi_{usc}$ at $v$ in the direction of $p$ equals $D\Phi_{usc}(v)(p)=-q$. Since $-D\Phi_{usc}(v):C_{usc}\to D_{usc}$ is a gauge-preserving bijection by Theorem 4.7, for each $g\in C$ we have that
$M(q,{-D\Phi}(v)(g))=M({-D\Phi_{usc}}(v)(p),{-D\Phi_{usc}}(v)(g))=M(p,g).$
Moreover, Proposition 3.18 grants us a state $\psi$ on $(W,W_{+},w)$ such that $M(p,g)=\Phi(g)(\psi)$ for each $g\in C$. An application of these results to $g=\Phi^{-1}(w+\Phi(y))$ yields
$M(q,{-D\Phi(v)}(\Phi^{-1}(w+\Phi(y))))=M(p,\Phi^{-1}(w+\Phi(y)))=w(\psi)+\Phi(y)(\psi)=1+M(p,y).$
Because $L$ and $M$ are multiplicatively inverse, we have $m(y,p)=M(p,y)^{-1}$ and
$m(w-\Phi(v+y)/q)M(q,{-D\Phi(v)}(\Phi^{-1}(w+\Phi(y))))=\frac{m(y,p)}{m(y,p)+1}(1+M(p,y))=1.$
For the same reason this equality is equivalent to (55), whence the proof is complete. ∎
We obtain the following formula for the Gateaux derivative $D\Phi(x)$ of $\Phi$ at points $x\in C$ in directions $u\in C$.
Corollary 4.11.
For all $x,u\in C$ with $2u\leq x$, the Gateaux derivative of $\Phi$ at $x$ in the direction $u$ is given by
$D\Phi(x)(u)=\Phi(x+\Phi^{-1}(\Phi(u)-\Phi(x)))-\Phi(x).$ (57)
Proof.
Since $2u\leq x$ gives $\Phi(x)\leq\frac{1}{2}\Phi(u)$, we have $\Phi(u)-\Phi(x)\in D$. As a result, the element $y=\Phi^{-1}(\Phi(u)-\Phi(x))\in C$ is well-defined and satisfies $\Phi(x)+\Phi(y)=\Phi(u)$. Therefore, Hua’s identity (54) yields $D\Phi(x)(u)=\Phi(x+y)-\Phi(x)$, which is (57). ∎
5 Proof of the main theorem
In this section we perform the analysis required to prove that gauge-reversing bijections are analytic as Theorem 5.19. We proceed in four stages, consisting of showing first continuity, then differentiability, then twice differentiability and ultimately analiticity. We strengthen Gateaux derivative positive directions is Fréchet derivative. is analytic, hence infinitely differentiable. Then define quadratic rep, and show it is a quadratic map. This ultimately allows us to prove analyticity.
5.1 Differentiability
In Theorem 4.7 we have established that for each $x\in C$ there exists an isomorphism of order unit spaces $-D\Phi(x):(V,V_{+},x)\to(W,W_{+},\Phi(x))$ such that if $y\in V_{+}$ then $D\Phi(x)(y)$ is the Gateaux-derivative of $\Phi$ at $x$ in the direction $y$. In this section we show that $D\Phi(x)$ is in fact the Fréchet-derivative of $\Phi$ at $x$.
The space $L(V,W)$ of all bounded linear operators $A\colon V\to W$ is a Banach space in the operator norm
$\lVert A\rVert_{v,w}:=\sup\{\lVert Ax\rVert_{w}:x\in V,\lVert x\rVert_{v}\leq 1\}.$
Lemma 5.1.
For each $A\in L(V,W)$ we have
$\lVert A\rVert_{v,w}\leq 3\sup\{\lVert Ax\rVert_{w}:x\in C,\lVert x\rVert_{v}\leq 1\}.$ (58)
Proof.
Every $x\in V$ can be written as $x=x_{1}-x_{2}$ with $x_{1},x_{2}\in V_{+}$ and $\lVert x_{1}\rVert_{v}+\lVert x_{2}\rVert_{v}\leq 3\lVert x\rVert_{v}$. Indeed, we may take $x_{2}=\lVert x\rVert_{v}v$ and $x_{1}=x+x_{2}$. The lemma follows since $C\cap[0,v]$ is $\lVert\cdot\rVert_{v}$-dense in $[0,v]$. ∎
Lemma 5.2.
For all $x\in C$ we have
$\lVert D\Phi(x)\rVert_{v,w}\leq M(v,x)^{2}.$ (59)
Proof.
The isomorphism of order unit spaces $-D\Phi(x)\colon(V,x)\to(W,\Phi(x))$ is an isometry with respect to order unit norms, that is, $\lVert-D\Phi(x)y\rVert_{\Phi(x)}=\lVert y\rVert_{x}$ for each $y\in V$. The order unit norms $\lVert\cdot\rVert_{x}$ and $\lVert\cdot\rVert_{v}$ on $V$ satisfy $\lVert\cdot\rVert_{x}\leq M(v,x)\lVert\cdot\rVert_{v}$. Likewise the norms $\lVert\cdot\rVert_{w}$ and $\lVert\cdot\rVert_{\Phi(x)}$ on $W$ are related by $\lVert\cdot\rVert_{w}\leq M(\Phi(x),w)\lVert\cdot\rVert_{\Phi(x)}$. Since $\Phi$ is gauge-reversing, it follows that
$\lVert D\Phi(x)y\rVert_{w}\leq M(\Phi(x),w)\lVert D\Phi(x)y\rVert_{\Phi(x)}=M(v,x)\lVert y\rVert_{x}\leq M(v,x)^{2}\lVert y\rVert_{v}$
for every $y\in V$, whence $\lVert D\Phi(x)\rVert_{v,w}\leq M(v,x)^{2}.$
The following lemma will be used to show continuity of the map $D\Phi:C\to L(V,W)$.
Lemma 5.3.
Let $\lambda>0$. For all $x,y,u\in C$ with $x,y\geq\lambda^{-1}v\geq 2u$ we have
$\lVert(D\Phi(x)-D\Phi(y))(u)\rVert_{w}\leq 3\lambda^{2}\lVert x-y\rVert_{v}.$ (60)
Proof.
Corollary 4.11 applied twice to $x\geq 2u$ and $y\geq 2u$ gives
$D\Phi(x)(u)-D\Phi(y)(u)=\Phi(x+\Phi^{-1}(\Phi(u)-\Phi(x)))-\Phi(y+\Phi^{-1}(\Phi(u)-\Phi(y)))-\Phi(x)+\Phi(y).$ (61)
By our assumption $x,y\geq\lambda^{-1}v$, the Lipschitz property (18) of $\Phi$ gives that
$\lVert\Phi(x)-\Phi(y)\rVert\leq\lambda^{2}\lVert x-y\rVert_{v}.$ (62)
Because $\Phi(u)-\Phi(x)\geq\Phi((2\lambda)^{-1}v)-\Phi(\lambda^{-1}v)=2\lambda w-\lambda w=\lambda w$ and likewise $\Phi(u)-\Phi(y)\geq\lambda w$, we find again by (18) that
$\lVert\Phi^{-1}(\Phi(u)-\Phi(x))-\Phi^{-1}(\Phi(u)-\Phi(y))\rVert_{v}\leq\lambda^{-2}\lVert\Phi(x)-\Phi(y)\rVert_{w}\\ \leq\lVert x-y\rVert_{v}.$ (63)
Once more applying (18), this time to $x+\Phi^{-1}(\Phi(u)-\Phi(x))$ and $y+\Phi^{-1}(\Phi(u)-\Phi(y))$, gives using the triangle inequality and (63) that
$\lVert\Phi(x+\Phi^{-1}(\Phi(u)-\Phi(x)))-\Phi(y+\Phi^{-1}(\Phi(u)-\Phi(y)))\rVert_{w}\leq 2\lambda^{2}\lVert x-y\rVert_{v}.$ (64)
Using the triangle inequality on (61), it follows from (64) and (62) that
$\lVert D\Phi(x)(u)-D\Phi(y)(u)\|_{w}\leq 2\lambda^{2}\lVert x-y\rVert_{v}+\lambda^{2}\lVert x-y\rVert_{v}$ $=3\lambda^{2}\lVert x-y\rVert_{v}.$
Proposition 5.4.
Let $\lambda>0$. For all $x,y\in C$ with $x,y\geq\lambda^{-1}v$ one has
$\lVert D\Phi(x)-D\Phi(y)\rVert_{v,w}\leq 18\lambda^{3}\lVert x-y\rVert_{v}.$
Moreover, for all $u\in V_{+}$ one has
$\lVert(D\Phi(x)-D\Phi(y))(u)\rVert_{w}\leq 6\lambda^{3}\lVert x-y\rVert_{v}\lVert u\rVert_{v}.$ (65)
Proof.
In view of Lemma 5.1, it suffices to show the last estimate.
First assume that $u\in C$. Then $\hat{u}=(2\lambda\lVert u\rVert_{v})^{-1}u$ satisfies $2\hat{u}\leq\lambda^{-1}v$. By the result of the previous paragraph, and the linearity of $D\Phi(x)-D\Phi(y)$ we obtain
$\lVert(D\Phi(x)-D\Phi(y))u\rVert_{w}=2\lambda\lVert u\rVert_{v}\cdot\lVert D\Phi(x)\hat{u}-D\Phi(y)\hat{u}\rVert_{w}\leq 2\lambda\lVert u\rVert_{v}\cdot 3\lambda^{2}\lVert x-y\rVert_{v}=6\lambda^{3}\lVert x-y\rVert_{v}\lVert u\rVert_{v}.$
Since $C$ is $\lVert\cdot\rVert_{v}$-dense in $V_{+}$, the inequality holds for all $u\in V_{+}$. ∎
We conclude that $D\Phi\colon C\to L(V,W)$ is locally Lipschitz continuous. The following lemma will be used to show that $D\Phi$ is in fact the Fréchet derivative of $\Phi$.
Lemma 5.5.
Let $\lambda>0$. For all $x\in C$ and $z\in V$ with $x,x+z\geq\lambda^{-1}v$ one has
$\lVert\Phi(x+z)-\Phi(x)-D\Phi(x)(z)\rVert_{w}\leq 11\lambda^{3}\lVert z\rVert_{v}^{2}.$
Proof.
As in the proof of Lemma 5.1, we can write $z=z_{2}-z_{1}$ with $0\leq z_{i}\leq i\lVert z\rVert_{v}v$ for $i=1,2$. In Theorem 4.7 we have shown that $D\Phi(x)(z_{2})$ is the Gateaux-derivative of $\Phi$ at $x$ in the direction $z_{2}\in V_{+}$ as a consequence of the estimate
$\lVert\Phi(x+z_{2})-\Phi(x)-D\Phi(x)(z_{2})\rVert_{\Phi(x)}\leq\lVert z_{2}\rVert_{x}^{2}.$
In view of the fact that $M(\Phi(x),v)=M(v,x)\leq\lambda$, we have $\lVert\cdot\rVert_{w}\leq\lambda\lVert\cdot\rVert_{\Phi(x)}$ and $\lVert\cdot\rVert_{x}\leq\lambda\lVert\cdot\rVert_{v}$, so our estimate entails that
$\lVert\Phi(x+z_{2})-\Phi(x)-D\Phi(x)(z_{2})\rVert_{w}\leq\lambda^{3}\lVert z_{2}\|_{v}^{2}.$
Similarly, because $M(v,x+z)\leq\lambda$ and $D\Phi(x+z)(z_{1})$ is the Gateaux-derivative of $\Phi$ at $x+z$ in the direction $z_{1}\in V_{+}$, we obtain
$\lVert\Phi(x+z+z_{1})-\Phi(x+z)-D\Phi(x+z)(z_{1})\rVert\leq\lambda^{3}\lVert z_{1}\rVert_{v}^{2}.$
Moreover, by the bound (65) in Proposition 5.4 we have
$\lVert(D\Phi(x+z)-D\Phi(x))(z_{1})\rVert_{v}\leq 6\lambda^{3}\lVert z\rVert_{v}\lVert z_{1}\rVert_{v}.$
Using the triangle inequality on
$\Phi(x+z)-\Phi(x)-D\Phi(x)z$ $=\Phi(x+z_{2})-\Phi(x)-D\Phi(x)z_{2}-(D\Phi(x+z)-D\Phi(x))z_{1}$
$\phantom{=}-(\Phi(x+z+z_{1})-\Phi(x+z)-D\Phi(x+z)z_{1}),$
it follows that
$\lVert\Phi(x+z)-\Phi(x)-D\Phi(x)z\rVert_{v}\leq\lambda^{3}(\lVert z_{2}\rVert_{v}^{2}+6\lVert z\rVert_{v}\lVert z_{1}\rVert_{v}+\lVert z_{1}\rVert_{v}^{2})\leq\lambda^{3}(2^{2}+6+1)\lVert z\rVert_{v}^{2}=11\lambda^{3}\lVert z\rVert_{v}^{2}.$
Theorem 5.6.
The map $\Phi:C\to D$ is continuously differentiable.
Proof.
Let $x\in C$, and set $\lambda=2m(x,v)^{-1}$. If $z\in V$ is small enough so that $\lVert z\rVert_{v}\leq\lambda^{-1}$, then $x+z\geq m(x,v)v-\lambda^{-1}v=\lambda^{-1}v$. Consequently, we may apply Lemma 5.5 to obtain that $\lVert\Phi(x+z)-\Phi(x)-D\Phi(x)(z)\rVert_{w}\leq 11\lambda^{3}\lVert z\rVert_{v}^{2}$. We conclude that $D\Phi(x)$ is the Fréchet-derivative of $\Phi$ at $x$. Similarly, by Proposition 5.4 we have $\lVert D\Phi(x+z)-D\Phi(x)\rVert\leq 18\lambda^{3}\lVert z\rVert_{v}$, hence $D\Phi:C\to L(V,W)$ is continuous at $x$. Since $x\in C$ was arbitrary, the theorem is proved. ∎
5.2 Symmetries and the quadratic representation
In this subsection, we define and construct symmetries of $C$ at each of its points. As a particular case, the symmetry $j=S_{v}$ at $v$ will later feature as the inversion map of the JB-algebra structure on $V$ with identity element $v$. Moreover, we introduce the quadratic representation and prove various formulae for it.
The following lemma shows that a gauge-reversing bijection is determined by its linearization.
Lemma 5.7.
Let $\Phi,\Upsilon\colon C\to D$ be two gauge-reversing bijections. If there exists $x\in C$ such that $\Phi(x)=\Upsilon(x)$ and $D\Phi(x)=D\Upsilon(x)$, then $\Phi$ and $\Upsilon$ are equal.
Proof.
Since $\Phi$ and $\Upsilon^{-1}$ are gauge-reversing bijections, we deduce that the composition $T=\Upsilon^{-1}\circ\Phi$ is a gauge-preserving bijection, hence extends to a bounded linear map on $V$ by Theorem 4.9. Two applications of the chain rule, and our assumptions on $\Phi$ and $\Upsilon$ at $x$ give that
$DT(x)=D(\Upsilon^{-1})(\Phi(x))\circ D\Phi(x)=D(\Upsilon^{-1})(\Upsilon(x))\circ D\Upsilon(x)=\operatorname{id}_{V}.$
It follows that $T=DT(x)|_{C}=\operatorname{id}_{C}$, whence $\Phi=\Upsilon\circ T=\Upsilon$. ∎
Definition 5.8.
For every $x\in C$ we define a symmetry of $C$ at $x$ to be a gauge-reversing bijection $S_{x}\colon C\to C$ such that $S_{x}(x)=x$ and $DS_{x}(x)=-\operatorname{id}_{V}$.
By Lemma 5.7, if a symmetry $S_{x}$ of $C$ at $x$ exists, it is unique. Existence is granted by the next lemma.
Lemma 5.9.
Let $\Upsilon\colon C\to D$ be any gauge-reversing bijection. Then for every $x\in C$, the map $S_{x}\colon C\to C$ defined by
$S_{x}=-[D\Upsilon(x)]^{-1}\circ\Upsilon$
is the symmetry of $C$ at $x$.
Proof.
Recall from Theorem 4.7 that $-D\Upsilon(x)\colon(V,V_{+},x)\to(W,W_{+},\Upsilon(x))$ is an isomorphism of order unit spaces. It follows that $S_{x}:C\to C$ is a well-defined gauge-reversing bijection with $S_{x}(x)=x$. Since $DS_{x}(x)=-[D\Upsilon(x)]^{-1}\circ D\Upsilon(x)=-\operatorname{id}_{V}$ by the chain rule, we conclude that $S_{x}$ is a symmetry of $C$ at $x$. ∎
In Definition 2.26 we defined a $d_{T}$-symmetry at a point $x\in C$ to be an involutive $d_{T}$-isometry having $x$ as an isolated fixed point. According to the following lemma, the symmetries of $C$ are $d_{T}$-symmetries and in particular are involutory.
Lemma 5.10.
Let $x\in C$ and suppose $S_{x}\colon C\to C$ is the symmetry of $C$ at $x$. Then $S_{x}$ is a $d_{T}$-symmetry at $x$.
Proof.
Observe that $S_{x}^{2}$ is a gauge-preserving bijection, which extends to a linear automorphism of $V$ by Theorem 2.24. Since $D(S_{x}^{2})(x)=DS_{x}(x)\circ DS_{x}(x)=-\operatorname{Id}_{V}\circ-\operatorname{Id}_{V}=\operatorname{Id}_{V}$, we find that $S_{x}^{2}=\operatorname{Id}_{C}$. As in the proof of [LRW25, Theorem 7.4], it follows from $DS_{x}(x)=-\operatorname{Id}_{V}$ that $x$ is an isolated fixed point of $S_{x}$. ∎
For the remainder of this subsection, we assume that some gauge-reversing bijection $\Phi\colon C\to D$ exists. For $x\in C$, we denote by $S_{x}$ the unique symmetry of $C$ at $x$. From Lemma 5.10 we see that $S_{x}$ is involutory, i.e. satisfies $S_{x}^{2}=\operatorname{id}_{C}$. The gauge-preserving bijection $S_{x}\circ S_{v}:C\to C$ extends by Theorem 4.9 to an automorphism $P(x)$ of $(V,V_{+})$. It follows that $S_{x}=S_{x}\circ S_{v}\circ S_{v}=P(x)\circ S_{v}$, so any two symmetries are related by a linear automorphism of $(V,V_{+})$.
Definition 5.11.
Let $(V,V_{+},v)$ be a complete order unit space with open cone $C:=V_{+}^{\circ}$. Assume that for each $x\in C$ there exists a symmetry $S_{x}\colon C\to C$ of $C$ at $x$.
  1. (1)
    We denote $j:C\to C$ the symmetry $S_{v}$ of $C$ at the given order unit $v$.
  2. (2)
    We define the quadratic representation of an element $x\in C$ to be the automorphism $P(x)$ of $(V,V_{+})$ determined by $S_{x}=P(x)\circ j$.
Note that $P(x)|_{C}=S_{x}\circ j$, hence it follows from Lemma 5.9 that for all $x\in C$ we have
$P(x)=[-Dj(x)]^{-1}=-Dj(j(x)).$ (66)
Hua’s identity (54) for the symmetry $S_{x}$, which satisfies $S_{x}(x)=x$ and $-DS_{x}(x)=\operatorname{id}_{V}$, reads
$S_{x}(x+y)+S_{x}(x+S_{x}(y))=x.$ (67)
Corollary 5.12.
For all $x,y\in C$ we have
$P(x)(y)=j(j(x)-j(x+j(y)))-x$ (68)
Proof.
Since $S_{x}=P(x)\circ j$ gives that $x=S_{x}(x)=P(x)j(x)$, applying $P(x)^{-1}$ to (67) gives that
$j(x+y)+j(x+P(x)j(y))=j(x)$ (69)
We may replace $y$ by $j(y)$, and rewrite this as
$j(x+P(x)y)=j(x)-j(x+j(y)).$
Since $j^{2}=\operatorname{id}_{C}$, applying $j$ to both members gives
$x+P(x)y=j(j(x)-j(x+j(y))).$
Subtract $x$ to arrive at (68). ∎
Theorem 5.13.
For all $x,y\in C$ we have
$(P(x)-P(y))j(x+y)=x-y.$ (70)
Proof.
Recall from Definition 5.11 and Lemma 5.10 that $P(x)\circ j=S_{x}=S_{x}^{-1}=j\circ P(x)^{-1}$ and $S_{x}(x)=x$. From Hua’s identity (67) for the symmetries $S_{x}$ and $S_{y}$ we obtain that
$(P(x)-P(y))j(x+y)=S_{x}(x+y)-S_{y}(x+y)=x-y-(S_{x}(x+S_{x}y)-S_{y}(S_{y}x+y)).$
The final term cancels since
$S_{x}(x+S_{x}(y))=S_{x}^{-1}(S_{x}(x)+S_{x}(y))=(j\circ P(x)^{-1})(P(x)(j(x))+P(x)(j(y)))=j(j(x)+j(y)),$
and similarly with $x$ and $y$ interchanged. ∎
We record two consequences of the uniqueness of symmetries.
Lemma 5.14.
For all $x,y\in C$ we have $S_{x}S_{y}S_{x}=S_{S_{x}(y)}$.
Proof.
It suffices to show that $S_{x}S_{y}S_{x}$ is the unique symmetry of $C$ at $S_{x}(y)$. Note first that $(S_{x}S_{y}S_{x})(S_{x}(y))=(S_{x}S_{y})(y)=S_{x}(y)$. Since $DS_{y}(y)=-\operatorname{id}_{V}$, using the chain rule we obtain $D(S_{x}S_{y}S_{x})(S_{x}(y))=DS_{x}(y)\circ DS_{y}(y)\circ DS_{x}(S_{x}(y))=-DS_{x}(y)\circ DS_{x}(S_{x}(y))=-D(S^{2}_{x})(S_{x}(y))$. Since $S^{2}_{x}=\operatorname{id}_{C}$, it follows that $D(S_{x}S_{y}S_{x})(S_{x}(y))=-\operatorname{Id}_{V}$. In view of Definition 5.8, we conclude that the map $S_{x}S_{y}S_{x}$ is equal to the symmetry $S_{S_{x}(y)}$ of $C$ at $S_{x}(y)$. ∎
Lemma 5.15.
For all $x,y\in C$ we have
$P(x)P(y)P(x)=P(P(x)y).$ (71)
Proof.
By repeated applications of Definition 5.11 and Lemma 5.14 we find that
$P(x)P(y)P(x)=S_{x}(S_{v}S_{y}S_{v})S_{x}S_{v}=(S_{x}S_{S_{v}(y)}S_{x})S_{v}=S_{S_{x}(S_{v}(y))}S_{v}=S_{P(x)y}S_{v}=P(P(x)y).$
5.3 Twice differentiability
We have shown in Theorem 5.6 that each gauge-reversing bijection $\Phi\colon C\to D$ is continuously differentiable. The aim of this subsection is to show in Theorem 5.17 that $\Phi$ is even twice continuously differentiable.
Let us first show that the map $P\colon C\to L(V)$ is continuously differentiable. For every $u\in V$, let $\operatorname{ev}_{u}\colon L(V)\to V$ be the continuous linear map given by evaluation in $u$. Define $Q\colon C\times V\to L(V)$ by
$Q(x,u):=D(\operatorname{ev}_{u}\circ P)(x).$
It is plain that $Q$ is linear in its second variable. Note that if $u\in C$ then
$(\operatorname{ev}_{u}\circ P)(x)=P(x)(u)=j(j(x)-j(x+j(u)))-x,$
so by the chain rule and Fréchet differentiability of $j$ we have
$Q(x,u)=D(\operatorname{ev}_{u}\circ P)(x)=Dj(j(x)-j(x+j(u)))\circ(Dj(x)-Dj(x+j(u)))-\operatorname{Id}.$ (72)
Since $Dj\colon C\to L(V)$ is continuous by Proposition 5.4, it follows that $Q$ is continuous on $C\times C$, and then by linearity on all of $C\times V$. Hence there exists a continuous map $DP\colon C\to L(V,L(V))$ such that for all $x\in C$ and $z,u\in V$ one has
$DP(x)(z)(u)=Q(x,u)(z).$
Theorem 5.16.
Let $\lambda>1$. Then for all $x\in C$ and $z\in V$ such that $x,x+z\in[\lambda^{-1}v,\lambda v]$ and $\lVert z\rVert_{v}\leq(2\lambda)^{-3}$ we have
$\lVert P(x+z)-P(x)-DP(x)(z)\rVert_{v}\leq 10^{4}\lambda^{6}\lVert z\rVert_{v}^{2}.$
Proof.
Fix $u\in C$ with $\lVert u\rVert_{v}\leq\lambda^{-1}$. We shall exhibit a positive integer $N_{0}$ such that
$\lVert(P(x+z)-P(x)-DP(x)(z))u\rVert_{v}\leq N_{0}\lambda^{5}\lVert z\rVert_{v}^{2}.$ (73)
Set $y:=x+j(u)$ and define the quantities
$a(z)$ $:=j(x)-j(y)+(Dj(x)-Dj(y))z,$
$E_{0}(z)$ $:=j(x+z)-j(y+z)-a(z),$
$E_{1}(z)$ $:=j(j(x+z)-j(y+z))-j(a(z)),$
$E_{2}(z)$ $:=j(a(z))-j(j(x)-j(y))-Dj(j(x)-j(y))(Dj(x)-Dj(y))(z).$
By substitution of (68) and (72) we obtain
$(P(x+z)-P(x)-DP(x)(z))u$ $=$
$j(j(x+z)-j(x+z+j(u))-(x+z)-j(j(x)-j(x+j(u)))+x$
$-Dj(j(x)-j(y))(Dj(x)-Dj(y))(z)+z$ $=E_{1}(z)+E_{2}(z).$
Since $y\geq x\geq\lambda^{-1}v$ and $y+z\geq x+z\geq\lambda^{-1}v$ we have by Lemma 5.5 that
$\lVert j(x+z)-j(x)-Dj(x)z\rVert_{v}\leq 11\lambda^{3}\lVert z\rVert_{v}^{2}$
and
$\lVert j(y+z)-j(y)-Dj(y)z\rVert_{v}\leq 11\lambda^{3}\lVert z\rVert_{v}^{2}.$
It follows that
$\lVert E_{0}(z)\rVert_{v}\leq 2\cdot 11\lambda^{3}\lVert z\rVert_{v}^{2}.$
We have $u\leq\lambda^{-1}v\leq j(x)$, so using (68) we find, since $P(x)j(x)=S_{x}(x)=x$, that
$j(x)-j(y)=j(x+P(x)u)\geq j(x+P(x)j(x))=j(2x)\geq(2\lambda)^{-1}v.$
The same holds with $x$ replaced by $x+z$, hence
$j(x+z)-j(y+z)\geq(2\lambda)^{-1}v.$
Since $y\geq x\geq\lambda^{-1}v$, we have by Lemma 5.2 that $\lVert Dj(x)z\rVert_{v}\leq\lambda^{2}\lVert z\rVert_{v}$ and $\lVert Dj(y)z\rVert_{v}\leq\lambda^{2}\lVert z\rVert_{v}$, whence
$\lVert(Dj(x)-Dj(y))z\rVert_{v}\leq 2\lambda^{2}\lVert z\rVert_{v}\leq\frac{2\lambda^{2}}{8\lambda^{3}}=\frac{1}{4\lambda}.$
It follows that
$a(z)=j(x)-j(y)+(Dj(x)-Dj(y))z\geq(2\lambda)^{-1}v-(4\lambda)^{-1}v=(4\lambda)^{-1}v.$
Since $j(x+z)-j(y+z),a(z)\geq(4\lambda)^{-1}v$, it follows by (18) that
$\lVert E_{1}(z)\rVert_{v}\leq(4\lambda)^{2}\lVert E_{0}(z)\rVert_{v}\leq(4\lambda)^{2}\cdot 2\cdot 11\lambda^{3}\lVert z\rVert_{v}^{2}=:N_{1}\lambda^{5}\lVert z\rVert_{v}^{2}.$
Finally, since $a(z),j(x)-j(y)\geq(4\lambda)^{-1}v$, we obtain using Lemma 5.5 that
$\lVert E_{2}(z)\rVert_{v}\leq 11(4\lambda)^{3}\lVert(Dj(x)-Dj(y))z\rVert_{v}^{2}\leq 11(4\lambda)^{3}(2\lambda^{2}\lVert z\rVert_{v})^{2}=:N_{2}\lambda^{5}\lVert z\rVert_{v}^{2}.$
Since $(P(x+z)-P(x)-DP(x)(z))u=E_{1}(z)+E_{2}(z)$, we conclude that (73) holds with constant $N_{0}=N_{1}+N_{2}=4^{2}\cdot 2\cdot 11+11\cdot 4^{3}\cdot 2^{2}\leq 10^{4}/3$.
It now follows that for arbitrary $u\in C$ one has
$\lVert(P(x+z)-P(x)-DP(x)(z))u\rVert_{v}\leq N_{0}\lambda^{5}\lVert z\rVert_{v}^{2}\lVert u\rVert_{v}.$
We conclude using Lemma 5.1 that
$\lVert P(x+z)-P(x)-DP(x)(z)\rVert_{v}\leq 3N_{0}\lambda^{6}\lVert z\rVert_{v}^{2}\leq 10^{4}\lambda^{6}\lVert z\rVert_{v}^{2}.$
Theorem 5.17.
Each gauge-reversing bijection $\Phi\colon C\to D$ is twice continuously differentiable.
Proof.
Since $\Phi=-D\Phi(v)\circ j$, where $-D\Phi(v)\colon(V,\lVert\cdot\rVert_{v})\to(W,\lVert\cdot\rVert_{w})$ is an isometric isomorphism, it suffices to prove that $j$ is twice continuously differentiable. In Lemma 5.5 it has been shown that $j$ is continuously differentiable. Moreover, in Theorem 5.16 we showed that $P\colon C\to L(V)$ is continuously differentiable as well. Since $Dj=-P\circ j$, it follows that the map $Dj\colon C\to L(V)$ is continuously differentiable, that is, $j$ is twice continuously differentiable. ∎
5.4 Analyticity
Recall from Definition 5.11 that we defined the quadratic representation map $P\colon C\to L(V)$ to be the composition $P=-Dj\circ j$. Since $j$ is homogeneous of degree $-1$, the map $Dj\colon C\to L(V)$ is homogeneous of degree $-2$, and so $P=-Dj\circ j$ has degree 2×1=2-2$\times$-1=2- 2 × - 1 = 2. In the present section we extend $P$ to a quadratic polynomial map $P\colon V\to L(V)$, using the same derivation as in [McC77, Theorem 3.4]. As a consequence, we prove that gauge-reversing bijections are analytic.
We will now deduce from Theorem 5.13 and Theorem 5.17 that $P$ is a quadratic map
Theorem 5.18 (cf. [McC77, Theorem 3.4]).
The quadratic representation map $P\colon C\to L(V)$ extends to a continuous quadratic polynomial map $P\colon V\to L(V)$.
Proof.
First fix $z\in C$ and let $y:=j(z)\in C$. We will show that $\operatorname{ev}_{z}\circ P\colon C\to V$, $x\mapsto P(x)z$ extends to a quadratic polynomial map on $V$.
Replacing $x$ by $tx$ in Theorem 5.13 gives
$(t^{2}P(x)-P(y))j(tx+y)=tx-y.$
Differentiating with respect to $t$ gives
$2tP(x)j(tx+y)+(t^{2}P(x)-P(y))Dj(tx+y)(x)=x.$ (74)
Differentiate again at $t=0$ to find that
$2P(x)z=P(y)[D^{2}j(y)](x)(x).$ (75)
Since $j$ is twice continuously differentiable, its second differential $D^{2}j(y)\colon V\times V\to V$ at $y$ is a continuous symmetric bilinear map, see e.g. [Hör90, Theorem 1.1.8]. Therefore, formula (75) defines an extension of $\operatorname{ev}_{z}\circ P\colon C\to V$ to a quadratic polynomial map on $V$, as asserted. It follows from $V=C-C$ and linearity of $P(x)$ for each $x\in C$, that such an extension in fact exists for each $z\in V$. Hence we obtain a quadratic polynomial map $P\colon V\to\{T\colon V\to V\colon T\text{ linear}\}$.
The map $P$ being a quadratic polynomial, it satisfies for all $x,y\in V$ the parallelogram identity
$P(x+y)+P(x-y)=2P(x)+2P(y).$
By virtue of this identity, it follows from $V=C-C$ and continuity of $P(x)$ for each $x\in C$, that for arbitrary $x\in V$ the linear map $P(x)$ is continuous, i.e. $P(x)\in L(V)$. Similarly, from the continuity of $P$ on $C$ asserted by Theorem 5.17, one deduces that $P$ is continuous on all of $V=C-C$. Thus $P\colon V\to L(V)$ is a continuous quadratic polynomial map. ∎
Theorem 5.19.
Each gauge-reversing bijection $\Phi\colon C\to D$ is analytic.
Proof.
As in the proof of Theorem 5.17 it may be assumed that $\Phi=j$. For each $x\in C$ we have $-Dj(x)x=j(x)$, hence
$j(x)=P(x)^{-1}(x).$ (76)
Theorem 2.6 for the Banach algebra $L(V)$ asserts that $G=\{T\in L(V):T\text{ is invertible}\}$ is open and that $G\to G$, $T\mapsto T^{-1}$ is analytic. For each $x\in C$, we have $P(x)=-[Dj(x)]^{-1}\in G$. Moreover $P\colon V\to L(V)$ is analytic, because it is a continuous quadratic polynomial map by Theorem 5.18. Moreover, the evaluation map $L(V)\times V\to V$, $(T,x)\mapsto T(x)$ is continuous and bilinear, hence analytic. Because the composition of analytic maps is again analytic, it follows from (76) that $j$ is analytic. ∎
5.5 Verification of the quadratic Jordan algebra axioms
In this section we will prove that the quadratic map $P:V\to L(V)$ satisfies axioms (QJ1-QJ3) for a quadratic Jordan algebra, following [McC77].
Proposition 5.20.
The quadratic map $P$ satisfies (QJ1), i.e.
$P(v)=\operatorname{Id}_{V}.$ (77)
Proof.
Because $j$ is the symmetry of $C$ at $v$, we have that $Dj(v)=-\operatorname{Id}_{V}$. Therefore, it holds that $P(v):=-[Dj(v)]^{-1}=\operatorname{Id}_{V}$. ∎
Let us define the bilinear map $P\colon V\times V\to L(V)$ by
$P(x,z):=\tfrac{1}{2}(P(x+z)-P(x)-P(z)).$ (78)
Note that $P(x)=P(x,x)$.
Lemma 5.21.
For all $x\in C$ and $y\in V$ we have
$P(x,y)j(x)=y.$ (79)
Proof.
Differentiate the relation $P(x+ty)j(x+ty)=x+ty$ with respect to $t$ at $t=0$, to obtain $2P(x,y)j(x)+P(x)Dj(x)y=y$. Since $P(x)\circ Dj(x)=-[Dj(x)]^{-1}\circ Dj(x)=-\operatorname{Id}_{V}$ it follows that $2P(x,y)j(x)=2y$. ∎
Proposition 5.22.
The quadratic map $P$ satisfies (QJ2), i.e. for all $x,y\in V$ we have
$P(x)P(y)P(x)=P(P(x)y).$ (80)
Proof.
We have proved in Lemma 5.15 that the equation (80) holds if $x,y\in C$. In view of Theorem 5.18 this is a polynomial equation in $L(V)$ of degree $4$ in $x$ and of degree $2$ in $y$. Since $V=C-C$ it follows that (80) holds for all $x,y\in V$. ∎
On linearizing (80) we obtain for all $x,y,z\in V$ the identity
$P(x)P(y,z)P(x)=P(P(x)y,P(x)z).$ (81)
Proposition 5.23.
The quadratic map $P$ satisfies (QJ3), i.e. for all $x,y,z\in V$ we have
$P(x)P(y,z)x=P(P(x)y,x)z.$
Proof.
Since this is a polynomial equation in $V$ (of degree $3$ in $x$, and of degree $1$ in $y$ and in $z$), we may assume $x,y,z\in C$. Let $a,b\in V$. By (79) we have $P(a,x)j(x)=a$. Differentiate with respect $x$ in the direction $b$, to find $P(a,b)j(x)+P(a,x)Dj(x)(b)=0$, or
$P(a,b)j(x)=P(a,x)P(x)^{-1}(b).$ (82)
Substituting $a=P(x)y$ and $b=P(x)z$ gives, using (81) and (82), the desired identity
$P(x)P(y,z)x$ $=P(x)P(y,z)P(x)j(x)=P(P(x)y,P(x)z)j(x)=P(a,b)j(x)$
$=P(a,x)P(x)^{-1}(b)=P(P(x)y,x)P(x)^{-1}P(x)z$
$=P(P(x)y,x)z.$
Corollary 5.24.
The quadruple $(V,v,P)$ is a quadratic Jordan algebra.
Proof.
The map $P\colon V\to L(V)$ is a quadratic polynomial map by Theorem 5.18. It satisfies the quadratic Jordan algebra axioms (QJ1-3) by Proposition 5.20, Proposition 5.22 resp. Proposition 5.23. We conclude that the triple $(V,v,P)$ is a quadratic Jordan algebra in the sense of Definition 2.34. ∎
5.6 Verification of the quadratic JB-algebra axioms
In the previous section, we have shown that $(V,V_{+},v,P)$ is a quadratic Jordan algebra. To show this quadruple is a quadratic JB-algebra in the sense of Definition 2.41, it remains to be shown that $P(x)$ is positive with norm $\lVert P(x)\rVert_{v}=\lVert x\rVert_{v}^{2}$ for each $x\in V$.
Theorem 5.25.
The linear map $P(x)$ is positive for each $x\in V$.
Proof.
We have already seen for each $x\in C$ that $P(x)$ is an automorphism of $(V,V_{+})$ in Definition 5.11(2). Fix an arbitrary $x\in V$. Since $V_{+}$ is the closure of $C=j(C)$, it suffices to show that $P(x)(j(y))\in V_{+}$ for all $y\in C$. In view of (75) and the positivity of $P(y)$, we need only show that $[D^{2}j(y)](x)(x)\in V_{+}$. Owing to Kadison’s Theorem 2.11, it is equivalent to prove that $\psi([D^{2}j(y)](x)(x))\geq 0$ for each state $\psi\in K$.
So let $\psi\in K$, and define the function $f\colon I\to\mathbb{R}$ by $f(t):=\psi(j(y+tx))$ on the open interval $I:=\{t\in\mathbb{R}:y+tx\in C\}$ around 00. Because $\Phi$ is convex by Theorem 3.19 and $\psi$ is a positive functional, $f$ is convex as well. Since $\psi$ is a continuous linear functional and $j$ is twice continuously differentiable by Theorem 5.17, so is the map $f$, and $f^{\prime\prime}(0)=[D^{2}(\psi\circ j)(y)](x)(x)=\psi([D^{2}j(y)](x)(x))$. Since $f$ is convex, it follows that $\psi([D^{2}j(y)](x)(x))=f^{\prime\prime}(0)\geq 0$, as desired. ∎
The formula in the following theorem has been demonstrated in Theorem 5.13 for all $x,y\in C$, and will now be extended to its natural domain of definition.
Lemma 5.26.
For all $x,y\in V$ with $x+y\in C$ one has
$(P(x)-P(y))j(x+y)=x-y.$ (83)
Proof.
Let $U=\{(x,y)\in V\times V\colon x+y\in C\}$. Then $U$ is a connected open subset of the Banach space $(V^{2},\lVert\cdot\rVert_{(v,v)})$. The mapping $f\colon U\to V$, $f(x,y):=(P(x)-P(y))j(x+y)-(x-y)$ is analytic on $U$ by Theorem 5.19. By Theorem 5.13 we have $f(x,y)=0$ for all $(x,y)\in C\times C$. The identity principle for analytic functions (2.5) implies that $f$ vanishes identically on $U$. ∎
We have shown that the map $j$ is analytic in Theorem 5.19. We now show its power series expansion around $v$ is given by the familiar geometric series, following [McC77, Theorem 4.4].
Proposition 5.27.
There exists $0<r\leq 1$ such that for all $h\in V$ with $\lVert h\rVert_{v}<r$ one has
$j(v-h)=\sum_{k=0}^{\infty}h^{k}.$ (84)
Proof.
Since $j$ is analytic by Theorem 5.19, there exist $0<r\leq 1$ and for each $k\in\mathbb{Z}_{\geq 0}$ a continuous symmetric homogeneous polynomial map $c_{k}\colon V\to V$ such that for all $h\in V$ with $\lVert h\rVert<r$ one has
$j(v+h)=\sum_{k=0}^{\infty}(-1)^{k}c_{k}(h).$
We proceed to show for all $k\in\mathbb{Z}_{\geq 0}$ and $h\in V$ that
$c_{k}(h)=(-1)^{k}h^{k}.$
If $h=0$ then $c_{0}(0)=v=h^{0}$ and $c_{k}(0)=0=h^{k}$ for all $k\in\mathbb{Z}_{>0}$. Now assume $h\neq 0$. Let $t$ be a real number with $|t|<\lVert h\rVert_{v}^{-1}$. Then $v+th\in C$ holds, so Lemma 5.26 with $x=v$ and $y=th$ gives $(I-t^{2}P(h))j(v+th)=v-th.$ If in addition $t^{2}\lVert P(h)\rVert_{v}<1$, then $I-t^{2}P(h)$ is invertible in $L(V)$. The powers of $h$ in the Jordan algebra $V$ being given by $h^{2k}=P(h)^{k}v$ and $h^{2k+1}=P(h)^{k}h$, it follows that whenever $|t|<\min(\lVert h\rVert_{v}^{-1},\lVert P(h)\rVert_{v}^{-1/2})$ one has
$j(v+th)=(I-t^{2}P(h))^{-1}(v-th)=\left(\sum_{k=0}^{\infty}t^{2k}P(h)^{k}\right)(v-th)=\sum_{k=0}^{\infty}(-1)^{k}t^{k}h^{k}.$
Comparison of both series for $j(v+th)$ yields $c_{k}(h)=(-1)^{k}h^{k}$, whence (84). ∎
Remark 5.28.
We will shortly see in Proposition 5.31 that $\lVert P(x)\rVert_{v}=\lVert x\rVert_{v}^{2}$ for each $x\in V$, which implies that Proposition 5.27 holds with $r=1$.
Lemma 5.29.
For all $x\in V$ with $\lVert x\rVert_{v}<1$ we have
$v-x^{2}=2j(j(v-x)+j(v+x)).$
Proof.
Let $r>0$ be as in Proposition 5.27. Because $x\mapsto x^{2}=P(x)v$ is continuous by Theorem 5.18, there exists $0<r_{1}\leq r$ such that $\lVert x\rVert<r_{1}$ implies $\lVert x^{2}\rVert<r$ for each $x\in V$. Let $x\in V$ with $\lVert x\rVert<r_{1}$. Then for each $h\in\{x,-x,x^{2}\}$ we have $\lVert h\rVert<r$, hence the power series expansion (84) is valid. We obtain that
$j(v-x)+j(v+x)=\sum_{k=0}^{\infty}x^{k}+\sum_{k=0}^{\infty}(-1)^{k}x^{k}=2\sum_{k=0}^{\infty}x^{2k}=2j(v-x^{2}).$
Now apply $j$ to obtain $v-x^{2}=2j(j(v-x)+j(v+x))$ provided that $\lVert x\rVert<r_{1}$.
Both sides of this equation are analytic functions of $x$ on the open unit ball, and have just been shown to be equal on the open ball of radius $r_{1}>0$ around 00. In view of the identity principle, they are identically equal on the whole open unit ball. ∎
Lemma 5.30.
For all $x\in V$ with $-v\leq x\leq v$ we have $0\leq x^{2}\leq v$.
Proof.
First assume that $\lVert x\rVert_{v}<1$. By convexity of $j$ we have
$0\leq j(\tfrac{1}{2}j(v-x)+\tfrac{1}{2}j(v+x))\leq j(v)=v.$
Upon using Lemma 5.29, we obtain $0\leq v-x^{2}\leq v$. Now a limit argument shows this also holds if $\lVert x\rVert_{v}=1$. ∎
Proposition 5.31.
For each $x\in V$ we have $\lVert P(x)\rVert_{v}=\lVert x\rVert^{2}_{v}$.
Proof.
Since $P\in L(V)$ is positive, we have $\lVert P(x)\rVert_{v}=\lVert P(x)v\rVert_{v}=\lVert x^{2}\rVert_{v}.$ Thus our objective is to show that $\lVert x^{2}\rVert_{v}=\lVert x\rVert_{v}^{2}$ for each $x\in V$. Since both members of this equation are homogeneous of degree $2$ in $x$, it suffices to prove that $\lVert x^{2}\rVert_{v}\leq 1$ if and only if $\lVert x\rVert_{v}\leq 1$.
Suppose that $\lVert x\rVert_{v}\leq 1$, i.e. $-v\leq x\leq v$. Then Lemma 5.30 gives $0\leq x^{2}\leq v$, whence $\lVert x^{2}\rVert_{v}\leq 1$. Conversely suppose that $\lVert x^{2}\rVert_{v}\leq 1$, i.e. $-v\leq x^{2}\leq v$. Using that $y^{2}\in V_{+}$ for each $y\in V$, we find that
$-v\leq\tfrac{1}{2}((x+v)^{2}-x^{2}-v)=x=\tfrac{1}{2}(x^{2}+v-(x-v)^{2})\leq v,$
whence $\lVert x\rVert_{v}\leq 1$. This concludes the proof of the proposition. ∎
Corollary 5.32.
The quadruple $(V,V_{+},v,P)$ is a (quadratic) JB-algebra.
Proof.
In Corollary 5.24 we have seen that $(V,v,P)$ is a quadratic Jordan algebra. For $(V,V_{+},v,P)$ to be a quadratic JB-algebra in the sense of Definition 2.41, we need to verify in addition, for each $x\in V$, that $P(x)$ is a positive operator with norm $\lVert P(x)\rVert_{v}=\lVert x\rVert_{v}^{2}$, which has been done in Theorem 5.25 and Proposition 5.31. ∎
5.7 Proof of the main theorems
In this section we collect the main results of this paper concerning complete order unit spaces. In particular, we complete the proof of Theorem 1.1 and Theorem 1.2 announced in the introduction.
The if direction of Theorem 1.1 has been shown in Theorem 2.42; the only if direction is part (4) of the following theorem.
Theorem 5.33.
Let $(V,V_{+},v)$ be a complete order unit space with open cone $C:=V_{+}^{\circ}$. Let $\Phi\colon C\to D$ be a gauge-reversing bijection to the open cone $D:=W^{\circ}_{+}$ of a second complete order unit space $(W,W_{+},w)$.
  1. $(1)$
    The map $\Phi\colon C\to D$ is bianalytic.
  2. $(2)$
    The map $-D\Phi(v):(V,V_{+},v)\to(W,W_{+},w)$ is an isomorphism of order unit spaces.
  3. $(3)$
    The map $S_{v}:=-[D\Phi(v)]^{-1}\circ\Phi$ is the unique symmetry of $C$ at $v$.
  4. $(4)$
    There exists a unique JB-algebra structure on $V$ with cone of squares $V_{+}$ and unit element $v$.
  5. $(5)$
    Its set of invertible squares is $C$, and inversion on $C$ is given by the symmetry $S_{v}$ of $C$ at $v$.
Remark 5.34.
Wright-Youngson have shown in [WY78, Theorem 4] the uniqueness assertion of part (4), i.e. that on a complete order unit space $(V,V_{+},v)$ there exists at most one JB-algebra structure with cone of squares $V_{+}$ and unit element $v$. Equivalently, they showed that if $V$ and $W$ are JB-algebras, then each isomorphism between the associated order unit spaces $(V,V_{+},v)$ and $(W,W_{+},w)$ is in fact a JB-algebra isomorphism. Below we will give a different proof of the uniqueness of a JB-algebra structure on $(V,V_{+},v)$ based on the uniqueness of a symmetry $S_{v}$ of $C$ at $v$.
Proof.
The gauge-reversing bijections $\Phi$ and $\Phi^{-1}$ are analytic by Theorem 5.19, hence $\Phi$ is bianalytic. In particular, $\Phi$ is differentiable, so part (2) is contained in Theorem 4.7. Part (3) is shown in Lemma 5.7 and Lemma 5.9. To ease notation set $j=S_{v}$. In Theorem 5.18 we have seen that the map $P\colon C\to\operatorname{Aut}((V,V_{+}))$ sending $x$ to $P(x):=-Dj(j(x))$ extends to a quadratic polynomial map $P\colon V\to L(V)$. In Corollary 5.32 we have established that $(V,V_{+},v,P)$ is a (quadratic) JB-algebra, whence the existence assertion in (4). We have encountered part (5) in (76). It remains to prove the uniqueness assertion in (4).
Suppose that $P^{\prime}\colon V\to L(V)$ is another structure of quadratic JB-algebra on $(V,V_{+},v)$ and let $j^{\prime}\colon C\to C$ be the corresponding inversion map. Then $j$ and $j^{\prime}$ are both symmetries of $C$ at $v$, so Lemma 5.7 concerning the uniqueness of symmetries gives that $j=j^{\prime}$. Now in any JB-algebra $A$ with unit element $1$, we have for all $x\in A^{\circ}_{+}$ Hua’s identity $x^{2}=(x^{-1}-(x+1)^{-1})^{-1}$. Indeed, this identity may easily be verified in the associative unital JB-subalgebra of $A$ generated by $x$. Using this identity and the fact that $V=C-C$ one shows that $j=j^{\prime}$ implies $P=P^{\prime}$ (cf. the second paragraph of the proof of Theorem 5.39). This establishes uniqueness of a JB-algebra structure on $(V,V_{+},v)$ with identity element $v$ and cone of squares $V_{+}$. ∎
We proceed to prove Theorem 1.2. The following theorem has been shown in [LRW25, Theorem 8.1] for reflexive order unit spaces, and more generally, for complete order unit spaces which satisfy $C=C_{usc}$ and $V=\operatorname{span}\operatorname{ext}C$.
Theorem 5.35.
Let $(V,V_{+},v)$ be a complete order unit space and let $C=V_{+}^{\circ}$ be its open cone. Then the following are equivalent:
  1. (1)
    we have that $(C,d_{T})$ is a symmetric Banach–Finsler manifold;
  2. (2)
    there exists a $d_{T}$-symmetry $S_{p}:C\to C$ at some point $p\in C$;
  3. (3)
    there exists a gauge-reversing bijection $\Phi\colon C\to D$ to the open cone $D:=W^{\circ}_{+}$ of a complete order unit space $(W,W_{+}.w)$;
  4. (4)
    there exists a unital JB-algebra structure on $(V,\lVert\cdot\rVert_{v})$ with cone of squares $V_{+}$.
Proof.
(1) $\implies$ (2): The definition of a symmetric Banach–Finsler manifold entails that there exists a $d_{T}$-symmetry of $C$ at each of its points, in particular at $p$.
(2) $\implies$ (3): Each $d_{T}$-symmetry of $C$ is a gauge-reversing bijection by [LRW25, Thm. 3.5].
(3) $\implies$ (4): This is part (4) of Theorem 5.33.
(4) $\implies$ (1): Choose a JB-algebra structure on $V$. Denote $j\colon C\to C$ the inversion map and $v$ its multiplicative identity. By Theorem 2.42 we have that $j\colon C\to C$ is a gauge-reversing bijection. Let $p\in C$ be an arbitrary point. Then the map $S_{p}:=-[Dj(p)]^{-1}\circ p$ is a $d_{T}$-symmetry at $p$ by Lemma 5.9 and Lemma 5.10. We conclude that $(C,d_{T})$ is a symmetric Banach–Finsler manifold. ∎
5.8 Gauge-reversing bijections on non-complete order unit spaces
In this section we will prove Theorem 1.3, which asserts that two order unit spaces are necessarily isomorphic if there exists a gauge-reversing bijection between their open cones. For complete order unit spaces this has been shown in Theorem 4.7; in order to reduce to this case, we will extend a gauge-reversing bijection between possibly non-complete order unit spaces to their completions. Finally, we exploit this extension to deduce from Theorem 5.33 that a non-complete order unit space that admits a gauge-reversing bijection to a second order unit space is a norm dense subalgebra of a JB-algebra. Let $(V,V_{+},v)$ be a, possibly non-complete, order unit space. In Lemma 2.14 we constructed a complete order unit space $(\hat{V},\hat{V}_{+},v)$ such that $(\hat{V},\lVert\cdot\rVert_{v})$ is the completion of the normed space $(V,\lVert\cdot\rVert_{v})$.
Lemma 5.36.
Let $(V,V_{+},v)$ be an order unit space, and let $(\hat{V},\hat{V}_{+},v)$ be its completion. The inclusion map $i\colon(C,d_{T})\to(\hat{C},d_{T})$ is a completion of the metric $(C,d_{T})$.
Proof.
The map $i$ is an isometry, because the ordering on $V$ is the restriction of the ordering on $\hat{V}$ to $V$. Hence it suffices to show that $C$ is $d_{T}$-dense in $\hat{C}$. Indeed, let $x\in\hat{C}$ be given and pick $\epsilon>0$ such that $x\geq\epsilon v$. Let $(y_{n})_{n}$ be a sequence in $V$ such that $\lVert y_{n}-x\rVert_{v}\to 0$. Set $x_{n}:=y_{n}+\lVert y_{n}-x\rVert_{v}v$ for each $n$, and note that $x_{n}\geq x$. Consequently $x_{n}\geq\epsilon v$, so Lemma 2.18 gives that $d_{T}(x_{n},x)\leq\epsilon^{-1}\lVert x_{n}-x\rVert_{v}\leq 2\epsilon^{-1}\lVert y_{n}-x\rVert_{v}$. We conclude that $d_{T}(x_{n},x)\to 0$ as $n\rightarrow\infty$. Since $x\in\hat{C}$ was arbitrary, this shows that $C$ is $d_{T}$-dense in $\hat{C}$ as desired. ∎
Lemma 5.37.
Let $\Phi\colon C\to D$ be a bijective $d_{T}$-isometry between the open cones of order unit spaces $V$ resp. $W$. Then there is a unique bijective $d_{T}$-isometry $\hat{\Phi}\colon\hat{C}\to\hat{D}$ extending $\Phi$. If $\Phi$ is gauge-preversing (resp. gauge-reserving), then $\hat{\Phi}$ is also gauge-preversing (resp. gauge-reserving).
Proof.
By the universal property of the completion of a metric space, there is a unique bijective isometry $\hat{\Phi}\colon(\hat{C},d_{T})\to(\hat{D},d_{T})$ extending $\Phi$.
Suppose that $\Phi$ is order-reserving. Let $x,y\in\hat{C}$ be such that $x\leq y$. We shall show that $\hat{\Phi}(x)\geq\hat{\Phi}(y)$. As in the proof of Lemma 5.36 we can find sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ in $C$ which $d_{T}$-converge to $x$ resp. $y$, and satisfy $x_{n}\leq x\leq y\leq y_{n}$ for each $n$. Since $\Phi$ is order-reversing, we have $\Phi(x_{n})\geq\Phi(y_{n})$ for each $n$. Moreover, we have that $d_{T}(x_{n},x)=d_{T}(\Phi(x_{n}),\hat{\Phi}(x))\to 0$, so Lemma 2.19 gives that $\lVert\Phi(x_{n})-\hat{\Phi}(x)\rVert_{w}\to 0$ as $n\to\infty$. Similarly, we have $\lVert\Phi(y_{n})-\hat{\Phi}(y)\rVert_{w}\to 0$ as $n\to\infty$. Because the cone $\hat{V}_{+}$ is the closure of $V_{+}$, it follows that $\hat{\Phi}(x)\geq\hat{\Phi}(y)$, which establishes that $\hat{\Phi}$ is order-preserving. Similarly, one may show that $\hat{\Phi}$ is order-preversing whenever $\Phi$ is order-preversing.
Assume that $\lambda>0$ and $\mu>0$ are such that $\mu\Phi(x)=\Phi(\lambda x)$ for all $x\in C$. The map $m_{C,\lambda}\colon C\to C$ given by $m_{C,\lambda}(x):=\lambda x$ for each $x\in C$ is a bijective $d_{T}$-isometry. We may reformulate our assumption as $m_{D,\mu}\circ\Phi=\Phi\circ m_{C,\lambda}$. It is plain that $\widehat{m_{C,\lambda}}=m_{\hat{C},\lambda}$, whence $m_{\hat{D},\mu}\circ\hat{\Phi}=\hat{\Phi}\circ m_{\hat{C},\lambda}$. We conclude that $\mu\hat{\Phi}(x)=\hat{\Phi}(\lambda x)$ for all $x\in\hat{C}$. Applying this result with $\lambda>0$ arbitrary and $\mu=\lambda$ (resp. $\mu=\lambda^{-1}$) shows that $\hat{\Phi}$ is homogeneous of degree $1$ (resp. of degree $-1$) whenever $\Phi$ is.
We conclude using Proposition 2.22 and Proposition 2.23. ∎
Theorem 1.3 will now be demonstrated.
Theorem 5.38.
Let $\Phi\colon C\to D$ be a gauge-reversing bijection between the open cones of order unit spaces $(V,V_{+},v)$ resp. $(W,W_{+},w)$. Then these order unit spaces are isomorphic.
Proof.
Lemma 5.37 yields a gauge-reserving bijection $\hat{\Phi}\colon\hat{C}\to\hat{D}$ between the open cones of the completions that extends $\Phi$. Now Theorem 4.7 furnishes the isomorphism of order unit spaces
$-D\hat{\Phi}(v)\colon(\hat{V},\hat{V}_{+},v)\to(\hat{W},\hat{W}_{+},w),$
whose inverse is given by $-D(\hat{\Phi}^{-1})(w)\colon(\hat{W},\hat{W}_{+},w)\to(\hat{V},\hat{V}_{+},v).$
Let us show that $D\hat{\Phi}(v)[V]\subset W$. First, consider $z\in C$ with $z\leq\tfrac{1}{2}v$. Then we obtain that $\Phi(z)-w\geq\Phi(\tfrac{1}{2}v)-w=2w-w=w,$ so $\Phi(z)-w\in D$. Applying Theorem 4.10 with $x=v$ and $y=\Phi^{-1}(\Phi(z)-w)$ gives
$\Phi(v+y)-D\hat{\Phi}(v)(\Phi^{-1}(\Phi(v)+\Phi(y)))=\Phi(v).$
Since $\Phi(v)+\Phi(y)=\Phi(z)$ and $\hat{\Phi}$ extends $\Phi$, we obtain
$-D\hat{\Phi}(v)(z)=w-\Phi(v+y)\in D.$
Since $\{z\in C:z\leq\tfrac{1}{2}v\}$ spans $V$ and $-D\hat{\Phi}(v)$ is linear, it follows that $-D\hat{\Phi}(v)[V]\subset W$. Similarly one shows that $-D(\widehat{\Phi}^{-1})(w)[W]\subset V$. We conclude that $-D\widehat{\Phi}(v)$ restricts to an isomorphism of order unit spaces from $(V,V_{+},v)$ onto $(W,W_{+},w)$. ∎
Theorem 5.39.
Let $(V,V_{+},v)$ be an order unit space. Suppose there exists a gauge-reversing bijection $\Phi\colon C\to D$ to a second order unit space $(W,W_{+},w)$. Then there exists a unique structure of Jordan algebra of $V$ with identity element $v$, and $V_{+}$ equals the closure of $\{x^{2}:x\in V\}$ in $V$ such that $V$ embeds isometrically as a subalgebra of a JB-algebra.
Proof.
Let $(\hat{V},\hat{V}_{+},v)$ be the completion of the order unit space $(V,V_{+},v)$. According to Lemma 5.37 we can extend $\Phi$ to a gauge-reversing bijection $\hat{\Phi}\colon\hat{C}\to\hat{D}.$ Invoke Theorem 5.33 to obtain a unique JB-algebra structure on $\hat{V}$ with cone of squares $\{x^{2}:x\in\hat{V}\}=\hat{V}_{+}$ and unit element $v$. Let us show that $V$ is a Jordan subalgebra of $\hat{V}$, i.e. that $x,y\in V$ implies $x\circ y\in V$.
Suppose first that $x,y\in C$, so that $x+y\in C$ as well. Identity (68) gives that
$x^{2}=P(x)v=j(j(x)-j(x+v))\in V,$
and similarly $y^{2},(x+y)^{2}\in V$. It follows that $x\circ y=\tfrac{1}{2}((x+y)^{2}-x^{2}-y^{2})\in V.$ Since $V=C-C$ and the product $\circ$ is bilinear, we obtain that $x\circ y\in V$ for all $x,y\in V$. Therefore $V$ is a norm dense Jordan subalgebra of the JB-algebra $\hat{V}$ containing the identity element $v$.
Let us now show that the closure of $\{x^{2}:x\in V\}$ in $V$ is equal to $V_{+}$. Theorem 5.33 asserts that $\{x^{2}:x\in\hat{V}\}=\hat{V}_{+}$. By consequence, $\{x^{2}:x\in V\}\subset\{x^{2}:x\in\hat{V}\}\cap V=\hat{V}_{+}\cap V=V_{+}$. Since $\hat{V}_{+}$ is closed in $\hat{V}_{+}$, it follows that $V_{+}=V\cap\hat{V}_{+}$ it closed in $V$ and the closure of $\{x^{2}:x\in V\}$ in $V$ is a subset of $V_{+}$. Conversely, suppose that $y\in V_{+}$. Then there exists $x\in\hat{V}_{+}$ with $x^{2}=y$. Since $V$ is dense in $\hat{V}$, we can choose a sequence $(x_{n})_{n}$ in $V$ converging to $x$. Because the squaring map is continuous in any JB-algebra, it follows that $(x_{n}^{2})_{n}$ converges to $x^{2}=y$. Thus $y$ belongs to the closure of $\{x^{2}:x\in V\}$ in $V$, as desired. This concludes the existence part of the theorem.
We now turn to the uniqueness part. Let $B$ be a JB-algebra containing $V$ isometrically as a Jordan subalgebra such that $V_{+}$ equals the closure of $\{x^{2}:x\in V\}$ in $V$ and $v$ is an identity element of $V$. Because the embedding of $V$ into $B$ is isometric, we may and will identify the completion $\hat{V}$ of $(V,\lVert\cdot\rVert_{v})$ with the closure of $V$ in $B$. Since the multiplication of $B$ is continuous, one sees that $\hat{V}$ is a JB-subalgebra of $B$ with identity element $v$. For the same reason, we have that the set $\{x^{2}:x\in\hat{V}\}$, which is closed in $\hat{V}$ since $\hat{V}$ is a JB-algebra, is equal to the closure of $\{x^{2}:x\in V\}$ in $\hat{V}$. Now we have assumed $V_{+}$ to be the closure of $\{x^{2}:x\in V\}$ in $V$, and the closure of $V_{+}$ in $\hat{V}$ is given by $\hat{V}_{+}$, from where $\{x^{2}:x\in\hat{V}\}=\hat{V}_{+}$. According to Theorem 5.33 there is at most one structure of JB-algebra on $\hat{V}_{+}$ having cone of squares $\hat{V}_{+}$ and unit element $v$, and it follows that the structure of Jordan algebra on $V$ having the properties stated in the theorem is unique as well. ∎